### v1

# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

Abstract.
We investigate linear parabolic equations in divergence form with singular coefficients and non-smooth boundary data. When the diffusion, drift, or potential terms, as well as the initial or boundary conditions, are distributions rather than functions, classical and weak solution concepts become inadequate due to the ill-posedness of products involving distributions. To overcome this, we introduce a framework of very weak solutions based on regularization techniques and the theory of moderate nets. Existence of very weak solutions is established under minimal regularity assumptions. We further prove consistency with classical solutions when the data are smooth and demonstrate uniqueness via negligibility arguments. Finally, we present numerical computations that illustrate the robustness of the very weak solution framework in handling highly singular inputs, including delta-type potentials and distributional boundary traces.
Key words and phrases:
parabolic equation, divergence form, shift term, energy methods, Galerkin approximation, singular coefficients, regularisation, very weak solution.
2020 Mathematics Subject Classification:
35K20, 35D30, 35K67.
This research was funded by the Science Committee of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP23486342), by the FWO Research Grant G083525N: Evolutionary partial differential equations with strong singularities, and by the Methusalem programme of the Ghent University Special Research Fund (BOF) (Grant number 01M01021).
1. Introduction
In this paper, we study a linear parabolic equation in divergence form with a drift term and potential on a bounded one-dimensional domain. We focus on the setting where the coefficients and the initial-boundary data are allowed to be singular, in the sense of distributions.
More precisely, we consider the initial–boundary value problem
$\partial_{t}u(t,x)-\partial_{x}\left(a(t,x)\partial_{x}u(t,x)\right)+b(t,x)\partial_{x}u(t,x)+q(x)u(t,x)=f(t,x),$
$\quad(t,x)\in[0,T]\times(0,1),$ (1.1)
with initial condition
$u(0,x)=u_{0}(x),\quad x\in(0,1),$ (1.2)
and inhomogeneous Dirichlet boundary conditions
$u(t,0)=g_{0}(t),\quad u(t,1)=g_{1}(t),\quad t\in[0,T].$ (1.3)
Here, $a(t,x)$ is the (possibly singular) diffusion coefficient associated with the divergence term, $b(t,x)$ represents the drift field, $q(x)$ is a spatial potential, $f(t,x)$ is an external source term, and $u_{0}(x)$, $g_{0}(t)$, and $g_{1}(t)$ are the prescribed initial and boundary data, respectively.
The well-posedness of such parabolic problems with regular coefficients and boundary data is a classical subject in the theory of partial differential equations (see, for instance, [6, 7, 8, 13]). However, when the diffusion coefficient $a$, the drift term $b$, the potential $q$, the source term $f$, the initial data $u_{0}$, or the boundary data $g_{0},g_{1}$ are allowed to be non-regular (e.g., involving Dirac delta functions or other distributions), the standard functional frameworks break down. In such cases, the product of distributions becomes undefined, and well-posedness in the classical or even weak sense fails to hold.
From an analytical standpoint, the study of PDEs with distributional coefficients lies at the frontier of modern functional analysis and the theory of generalized functions. The classical theory, built upon Sobolev spaces and variational methods, becomes inadequate in the presence of distributional multiplicative terms due to Schwartz’s impossibility theorem [28].
Similar challenges arise in hyperbolic equations. To address them, the notion of very weak solutions was introduced in [19], enabling analysis of second-order equations with singular coefficients. This approach has since been developed in numerous works addressing various types of singularities and equations (see [1, 2, 3, 4, 5, 9, 10, 11, 12, 17, 18, 20, 22, 23, 29]). Recent works on elliptic, hyperbolic and parabolic equations for the Sturm-Liouville operator with singular coefficients [21, 24, 25, 26] have successfully employed separation of variables to construct very weak solutions.
In particular, the parabolic case presents additional challenges due to the time evolution aspect, which complicates the treatment of singularities in a dynamic context. The extension of the very weak solution framework to such settings is both analytically rich and methodologically nontrivial. In our approach, we employ regularisation techniques to treat singular coefficients and data, and we derive uniform a priori estimates to control the behaviour of solution sequences. Moreover, we establish consistency with classical solutions in the smooth-data regime, ensuring the robustness and completeness of our methodology.
Singularities in model coefficients and data arise naturally in many physical and engineering applications. For instance, diffusion processes in heterogeneous media often involve abrupt changes or localised inclusions, which are well-modelled by delta functions or piecewise constant coefficients. In heat conduction, materials with phase boundaries or layered composites introduce sharp discontinuities in thermal properties. Similarly, in quantum mechanics, the inclusion of point interactions or singular potentials is essential in describing certain scattering phenomena. Fluid flow through porous media or semiconductor modelling may involve drift terms with abrupt transitions due to electric fields or impurities. The drift-diffusion equation with discontinuous or singular terms has been particularly significant in models of charged particle transport.
In these settings, the idealised or experimentally derived data often lack the regularity required by classical PDE theory. Hence, a generalised framework that accommodates singular coefficients is not only mathematically meaningful but also essential for the faithful representation of real-world phenomena.
From the computational perspective, the development of a rigorous analytical foundation for singular models has direct implications for numerical methods. Singularities are frequently encountered in discretised models, either due to coarse resolution or because of inherent features of the problem (e.g., sharp interfaces, point sources, or localised forcing). Without a sound analytical underpinning, numerical schemes may fail to converge, or worse, produce misleading results.
Our results provide justification for numerical approximations by showing that very weak solutions are limits of regularized problems for which standard numerical schemes (e.g., finite differences or finite elements) are applicable.
The purpose of this work is to extend the framework of very weak solutions to the linear parabolic equation (1)–(1.3) in the presence of both spatial and time-dependent singularities. Our analysis accommodates singularities in the diffusion coefficient $a$, drift term $b$, potential $q$, source term $f$, initial data $u_{0}$, and boundary data $g_{0},g_{1}$, by regularising the problematic terms and constructing solution sequences with uniform estimates. This strategy allows us to prove the existence, uniqueness, and consistency of very weak solutions with classical solutions when the data are smooth.
In a recent work by the authors [27], the analysis of the Klein–Gordon equation with singular coefficients was carried out under inhomogeneous Dirichlet boundary conditions. The analytical techniques developed in that study are applicable to the present problem, with appropriate modifications.
In this work, we extend the analytical framework to parabolic partial differential equations with singular coefficients and inhomogeneous Dirichlet boundary conditions in the one-dimensional spatial setting. Our formulation accommodates singularities in all components of the model—including diffusion, drift, potential, source terms, and boundary data—through a regularization strategy. We establish well-posedness in the very weak sense under this framework. Additionally, we present numerical simulations that illustrate the behaviour of the solutions. Our methodology builds upon and adapts analytical techniques previously developed for the Klein–Gordon equation with singularities [27].
The primary novelty of this work lies in the analysis of a parabolic equation with singular coefficients and a source term, under inhomogeneous Dirichlet boundary conditions — a setting which, to the best of our knowledge, has not yet been investigated in the literature. This study provides a robust theoretical framework for future analytical and numerical developments in this direction.
2. Homogeneous Dirichlet boundary condition
We now consider a linear parabolic equation in divergence form with drift and potential terms, with initial data and homogeneous Dirichlet condition:
$\left\{\begin{array}[]{l}\partial_{t}u-\partial_{x}(a(t,x)\partial_{x}u)+b(t,x)\partial_{x}u+q(x)u=f(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ u(0,x)=u_{0}(x),\quad x\in(0,1),\\ u(t,0)=u(t,1)=0\quad t\in[0,T],\end{array}\right$. (2.1)
where the diffusion and drift coefficients, respectively $a,b\in L^{\infty}([0,T];L^{\infty}(0,1))$ with $a(t,x)\geq\alpha>0$, the potential $q\in L^{\infty}(0,1)$, the source term $f\in L^{2}\big([0,T];L^{2}(0,1)\big)$, and the initial data $u_{0}\in L^{2}(0,1)$.
To build a weak solution of the problem (2.1) we use Galerkin’s method, as in [27].
We are looking for a weak solution of the problem (2.1) in the following sense:
Definition 1 (Weak solution).
A function $u\in L^{2}(0,T;H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ is said to be a weak solution to the problem (2.1), if the following identity holds for all test functions $\varphi\in H^{1}_{0}(0,1)$ and almost every $t\in[0,T]$:
$\frac{d}{dt}\int_{0}^{1}u(t,x)\varphi(x)\,dx$ $+\int_{0}^{1}a(t,x)\,\partial_{x}u(t,x)\,\partial_{x}\varphi(x)\,dx+\int_{0}^{1}b(t,x)\,\partial_{x}u(t,x)\,\varphi(x)\,dx$
$+\int_{0}^{1}q(x)\,u(t,x)\,\varphi(x)\,dx=\int_{0}^{1}f(t,x)\,\varphi(x)\,dx.$ (2.2)
2.1. Galerkin Approximation
We apply the Galerkin method to construct approximate solutions to the problem on the domain $0,T]\times(0,1)$. Let $\{w_{k}\}_{k=1}^{\infty}$ be a sequence of smooth functions satisfying
$\{w_{k}\}_{k=1}^{\infty}$ is an orthogonal basis of $H^{1}_{0}(0,1)$ and an orthonormal basis of $L^{2}(0,1)$. (2.3)
For each $m\in\mathbb{N}$, we define the Galerkin approximation
$u_{m}(t,x):=\sum_{k=1}^{m}d_{m}^{k}(t)\,w_{k}(x),$ (2.4)
with coefficients $\{d_{m}^{k}(t)\}_{k=1}^{m}$ determined by
$d_{m}^{k}(0)=(u_{0},w_{k})_{L^{2}(0,1)},$ (2.5)
and by requiring that $u_{m}$ satisfies the weak form projected onto each basis function:
$(\partial_{t}u_{m},w_{k})+(a\,\partial_{x}u_{m},\partial_{x}w_{k})+(b\,\partial_{x}u_{m},w_{k})+(qu_{m},w_{k})=(f,w_{k}),$ (2.6)
for all $k=1,\dots,m$ and $t\in[0,T]$.
Proposition 2.1.
For every $m\in\mathbb{N}$, there exists a unique function $u_{m}$ of the form (2.4) satisfying (2.5)–(2.6).
Proof.
Substituting (2.4) into (2.6) and using the orthonormality of $\{w_{k}\}$ in $L^{2}(0,1)$ yields
${d_{m}^{k}}^{\prime}(t)+\sum_{l=1}^{m}(a\,\partial_{x}w_{l},\partial_{x}w_{k})\,d_{m}^{l}(t)+\sum_{l=1}^{m}(b\,\partial_{x}w_{l},w_{k})\,d_{m}^{l}(t)+\sum_{l=1}^{m}(qw_{l},w_{k})\,d_{m}^{l}(t)$
$=(f(t),w_{k}),$
for $k=1,\dots,m$ and $t\in[0,T]$.
Let us define the coefficient functions:
$a^{lk}(t):=(a(t,\cdot)\,\partial_{x}w_{l},\partial_{x}w_{k}),\quad b^{lk}(t):=(b(t,\cdot)\,\partial_{x}w_{l},w_{k}),\quad g^{lk}:=(qw_{l},w_{k}),$
$f^{k}(t):=(f(t,\cdot),w_{k}).$
Then the resulting ODE system takes the form:
${d_{m}^{k}}^{\prime}(t)+\sum_{l=1}^{m}\left(a^{lk}(t)+b^{lk}(t)+g^{lk}\right)d_{m}^{l}(t)=f^{k}(t),$ (2.7)
with initial data $d_{m}^{k}(0)=(u_{0},w_{k})$. This is a linear system of ODEs for the vector $\mathbf{d}_{m}(t)=(d_{m}^{1}(t),\ldots,d_{m}^{m}(t))^{\top}$. Since all coefficients are measurable and bounded on $[0,T]$, standard ODE theory ensures the existence and uniqueness of a solution $\mathbf{d}_{m}\in C^{1}([0,T];\mathbb{R}^{m})$, from which the function $u_{m}$ follows via (2.4). ∎
2.2. Energy Estimates
We now derive uniform energy bounds for the Galerkin approximations $u_{m}$ as $m\to\infty$.
Theorem 2.2.
Let $u_{m}$ be the Galerkin approximation defined in (2.4), and assume that the coefficients satisfy
$a(t,x)\geq\alpha>0,\quad a,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$ (2.8)
Then, for all $m\in\mathbb{N}$, the following energy estimate holds:
$\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{t}u_{m}\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
[…]
$\qquad\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.9)
where the constant $C>0$ depends only on the final time $T$ and the spatial domain $(0,1)$, and is independent of $m$, the coefficients $a,b,q$, and the data $u_{0}$ and $f$.
Proof.
We test the weak form (2.6) with $u_{m}$. This yields
$Re(\partial_{t}u_{m},u_{m})+Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})+Re(b\,\partial_{x}u_{m},u_{m})+Re(qu_{m},u_{m})=Re(f,u_{m}).$
The first term is
$Re(\partial_{t}u_{m},u_{m})=\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
Using $a\geq\alpha>0$, the second term satisfies
$Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})\geq\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}.$
For the third term, integration by parts gives
$Re(b\,\partial_{x}u_{m},u_{m})=-\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}),$
so
$|(b\,\partial_{x}u_{m},u_{m})|\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
The potential term is
$Re(qu_{m},u_{m})=Re(\sqrt{q}u_{m},\sqrt{q}u_{m})=\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}.$
And the right-hand side is bounded by
[…]
Putting all together, we obtain
$\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}+2\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+2\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}\leq\|f\|^{2}_{L^{2}(0,1)}+C_{1}\|u_{m}\|^{2}_{L^{2}(0,1)},$
where $C_{1}:=1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}$.
Letting $\eta(t):=\|u_{m}(t)\|_{L^{2}(0,1)}^{2}$, this gives:
$\eta^{\prime}(t)\leq C_{1}\eta(t)+\|f(t)\|_{L^{2}(0,1)}^{2}.$
Applying Grönwall’s inequality, we obtain:
$\|u_{m}(t)\|_{L^{2}(0,1)}^{2}\leq e^{C_{1}t}\left(\|u_{0}\|_{L^{2}(0,1)}^{2}+\int_{0}^{t}\|f(s)\|_{L^{2}(0,1)}^{2}\,ds\right).$
Taking the supremum over $t\in[0,T]$, we deduce:
[…]
where $C$ depends only on $T$.
Integrating the energy inequality over $[0,T]$, and using the previous bound, we obtain:
$\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$
$\qquad\leq C\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.10)
where the constant $C$ depends only on $T$.
To estimate $\partial_{t}u_{m}$, let $\phi\in H^{1}_{0}(0,1)$ with $\|\phi\|_{H^{1}_{0}(0,1)}\leq 1$. Then, from the weak formulation, we have
$Re\langle\partial_{t}u_{m}(t),\phi\rangle$ $=Re(f(t),\phi)-Re(a(t)\,\partial_{x}u_{m}(t),\partial_{x}\phi)$
$-Re(b(t)\,\partial_{x}u_{m}(t),\phi)-Re(q\,u_{m}(t),\phi).$
Applying Hölder’s inequality and the Poincaré inequality on $(0,1)$, we estimate
$|\langle\partial_{t}u_{m}(t),\phi\rangle|\leq\|f(t)\|_{L^{2}(0,1)}\|\phi\|_{L^{2}(0,1)}+\|a(t)\|_{L^{\infty}(0,1)}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}\|\partial_{x}\phi\|_{L^{2}(0,1)}$
$\qquad+\|b(t)\|_{L^{\infty}(0,1)}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}\|\phi\|_{L^{2}(0,1)}+\|q\|_{L^{\infty}(0,1)}\|u_{m}(t)\|_{L^{2}(0,1)}\|\phi\|_{L^{2}(0,1)}$
$\qquad\leq C_{P}\left[\|f(t)\|_{L^{2}(0,1)}+\|b(t)\|_{L^{\infty}(0,1)}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}+\|q\|_{L^{\infty}(0,1)}\|u_{m}(t)\|_{L^{2}(0,1)}\right]$
$\qquad+\|a(t)\|_{L^{\infty}(0,1)}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)},$
where $C_{P}>0$ is the Poincaré constant for $(0,1)$. Taking the supremum over all $\phi\in H^{1}_{0}(0,1)$ with $\|\phi\|_{H^{1}_{0}(0,1)}\leq 1$, we obtain
$\|\partial_{t}u_{m}(t)\|_{H^{-1}(0,1)}$ $\leq C_{2}\Big(\|f(t)\|_{L^{2}(0,1)}+\left(\|a(t)\|_{L^{\infty}(0,1)}+\|b(t)\|_{L^{\infty}(0,1)}\right)\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}$
$+\|q\|^{\frac{1}{2}}_{L^{\infty}}\|\sqrt{q}u_{m}(t)\|_{L^{2}(0,1)}\Big),$
where
$C_{2}:=\max\left\{C_{P},1\right\}.$
Applying (2.2) for $u_{m}$ and $\partial_{x}u_{m}$ by integrating with respect to $t$ over $[0,T]$, we conclude
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;H^{-1}(0,1))}$ $\leq C\cdot\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}$
[…]
$\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\right),$ (2.11)
where $C$ depends only on $T$ and the domain $(0,1)$. This completes the proof of Theorem 2.2.
Theorem 2.3.
Let $T>0$, and consider the interval $(0,1)\subset\mathbb{R}$. Then, the problem (2.1) admits a unique weak solution
$u\in L^{2}([0,T];H^{1}_{0}(0,1))\,\,\,\text{with}\,\,\,\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1)),\,\,\,\sqrt{q}u\in L^{2}([0,T];L^{2}(0,1)),$
in the sense of Definition 1.
Proof.
Let $\{u_{m}\}_{m=1}^{\infty}$ be the Galerkin approximations constructed in (2.4). The uniform energy estimate (2.2) ensures that the following sequences are bounded independently of $m$:
$\{u_{m}\}\subset L^{2}([0,T];H^{1}_{0}(0,1)),\quad\{\partial_{t}u_{m}\}\subset L^{2}([0,T];H^{-1}(0,1)),$
$\{\sqrt{q}u_{m}\}\subset L^{2}([0,T];L^{2}(0,1)),\quad\{u_{m}\}\subset L^{\infty}([0,T];L^{2}(0,1)).$
By Banach–Alaoglu and the Aubin–Lions compactness lemma (see [15, Theorem 8, p. 289]), there exists a subsequence $u_{m_{l}}$ and a limit function $u$ such that:
$\left\{\begin{aligned} u_{m_{l}}&\rightharpoonup u&&\text{in }L^{2}([0,T];H^{1}_{0}(0,1)),\\ \sqrt{q}u_{m_{l}}&\rightharpoonup\sqrt{q}u&&\text{in }L^{2}([0,T];L^{2}(0,1)),\\ \partial_{t}u_{m_{l}}&\rightharpoonup\partial_{t}u&&\text{in }L^{2}([0,T];H^{-1}(0,1)),\\ u_{m_{l}}&\to u&&\text{in }L^{2}([0,T];L^{2}(0,1)).\end{aligned}\right$. (2.12)
Using these convergence properties and the weak continuity of all terms in the variational formulation, we can pass to the limit in the Galerkin formulation. Hence, $u$ satisfies the weak form of (2.1).
Uniqueness. Suppose $u\in L^{2}([0,T];H^{1}_{0}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$, and that $u$ is a weak solution corresponding to zero initial data and zero forcing term, i.e., $u_{0}=0$ and $f=0$. Then, for a.e. $t\in[0,T]$, testing the weak formulation with $u(t)$ yields:
$Re\langle\partial_{t}u(t),u(t)\rangle+Re(a\,\partial_{x}u(t),\partial_{x}u(t))+Re(b\,\partial_{x}u(t),u(t))+Re(qu(t),u(t))=0.$
Using the identity $Re\langle\partial_{t}u(t),u(t)\rangle=\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}$, and integrating by parts in the convective term:
$Re(b\,\partial_{x}u,u)=-\frac{1}{2}(\partial_{x}b\,,|u|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u\|^{2}_{L^{2}(0,1)}.$
Together with the assumption $a\geq\alpha>0$ and the positivity of $q$, we obtain the differential inequality:
$\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}+\alpha\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}+\|\sqrt{q}u(t)\|^{2}_{L^{2}(0,1)}\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u(t)\|^{2}_{L^{2}(0,1)}.$
Applying Grönwall’s inequality and using $u(0)=0$, we conclude that $u(t)\equiv 0$ for all $t\in[0,T]$, proving uniqueness. ∎
Proposition 2.4.
Let $u\in L^{2}([0,T];H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ be the weak solution to the parabolic problem (2.1), and assume the coefficients satisfy:
$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
Then, the weak solution to the parabolic equation (2.1) we have the following estimates:
$\Big\{\alpha^{1/2}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))};\left.\;\|u\|_{L^{\infty}([0,T];L^{2}(0,1))};\;\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}\right\}$
$\qquad\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.13)
$\|\partial_{t}u\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
[…]
$\qquad\times\left(\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|u_{0}\|_{L^{2}(0,1)}\right),$ (2.14)
$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2})};\,\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2})};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})};\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2})}\Big\}$
$\quad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\quad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.15)
where the constant $C>0$ depends only on the final time $T$ and the spatial interval $(0,1)$, but not on the data or coefficients.
Proof.
By taking limits in (2.2)-(2.2) as $m=m_{l}\to\infty$, we derive the estimates (2.4)-(2.4).
To prove (2.4), we start from the equation in strong form:
$\partial_{t}u-\partial_{x}(a\partial_{x}u)=-b\partial_{x}u-qu+f.$
Multiply both sides by $\partial_{t}u$ and integrate over $(0,1)$:
$\|\partial_{t}u\|^{2}_{L^{2}}$ $-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)+Re(b\partial_{x}u\,,\partial_{t}u)$
$+Re(qu\,,\partial_{t}u)=Re(f\,,\partial_{t}u).$
We handle the second term by integration by parts:
$-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)=Re(a\partial_{x}u\,,\partial_{x}(\partial_{t}u))=\frac{1}{2}\frac{d}{dt}(a\,,|\partial_{x}u|^{2})-\frac{1}{2}(\partial_{t}a\,,|\partial_{x}u|^{2})$
$\qquad\geq\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}-\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)},$
where the second term we move to the right hand-side.
Apply Cauchy–Schwarz and Young’s inequality to the remaining terms:
$\left|(b\partial_{x}u\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}},$
$\left|(qu\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}},$
$\left|(f\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
Combining everything, we obtain:
$\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}$ $\leq\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}$
$+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
Now integrate over $[0,T]$:
$\frac{1}{4}\|\partial_{t}u(t)\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u\|^{2}_{L^{\infty}([0,T];L^{2}(0,1))}$
$\qquad\leq\left(\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|^{2}_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\|\partial_{x}u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|q\|^{2}_{L^{\infty}(0,1)}\|u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\|f\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u(0)\|^{2}_{L^{2}(0,1)}.$
We now estimate the last term $\|\partial_{x}u(0)\|_{L^{2}}$ by using initial condition, we obtain
$\|\partial_{x}u(0)\|_{L^{2}(0,1)}=\|\partial_{x}u_{0}\|_{L^{2}(0,1)}=\|u_{0}\|_{H^{1}_{0}(0,1)}.$
Using the full energy estimate, we also have bounds for $\|u\|_{L^{2}([0,T];L^{2})}$, $\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}$ in terms of data. Therefore:
$\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))}+\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
$\qquad\times\left(\|a\|^{1/2}_{L^{\infty}}+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|_{L^{\infty}}\right)$
$\qquad\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|^{\frac{1}{2}}_{L^{\infty}}\right)\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$
where $C$ depends only on $T$ and the domain $(0,1)$.
Now, use the PDE again to write:
$\partial_{x}(a\partial_{x}u)=\partial_{t}u+b\partial_{x}u+qu-f.$
Then
$\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})}$ $\leq\|\partial_{t}u\|_{L^{2}}+\|b\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}}+\|q\|_{L^{\infty}}\|u\|_{L^{2}}+\|f\|_{L^{2}},$
which again is controlled by data via the energy estimate.
Finally, to estimate the second spatial derivative $\partial_{x}^{2}u$, we use the identity
$a\,\partial_{x}^{2}u=\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u,$
and the ellipticity condition $a(t,x)\geq\alpha>0$. Then, by the triangle inequality and Hölder’s inequality:
$\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$ $=\left\|\frac{1}{a}\left(\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u\right)\right\|_{L^{2}}$
$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\,\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right)$
$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\|_{L^{\infty}([0,T]\times(0,1))}\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right).$
Therefore, $\partial_{x}^{2}u\in L^{2}([0,T];L^{2}(0,1))$ with a constant depending only on $T$ and the interval $(0,1)$.
This completes the proof, with all terms estimated in terms of data and coefficients, and the constant depending only on $T$ and the interval $(0,1)$. ∎
3. Non-homogeneous Dirichlet condition
We now consider the parabolic equation in divergence form (1) with diffusion coefficient, drift term, and potential with initial condition (1.2) and non-homogeneous Dirichlet boundary conditions (1.3).
We decompose the solution as
$u(t,x)=w(t,x)+\psi(t,x),$ (3.1)
where $w$ satisfies homogeneous Dirichlet boundary conditions
$w(t,0)=w(t,1)=0,$
and $\psi$ is a lifting function defined by
$\psi(t,x):=(1-x)g_{0}(t)+xg_{1}(t),\quad(t,x)\in[0,T]\times(0,1).$
This function satisfies:
  • $\psi(t,0)=g_{0}(t),\quad\psi(t,1)=g_{1}(t)$,
  • $\partial_{x}\psi(t,x)=g_{1}(t)-g_{0}(t)$ is independent of $x$,
  • $\psi\in H^{1}([0,T];H^{1}(0,1))$ provided $g_{0},g_{1}\in H^{1}([0,T])$.
Substituting (3.1) into (1), we obtain the equation for $w$:
$\left\{\begin{array}[]{l}\partial_{t}w-\partial_{x}(a(t,x)\partial_{x}w)+b(t,x)\partial_{x}w+q(x)w=\tilde{f}(t,x),\\[2.0pt] w(0,x)=u_{0}(x)-\psi(0,x),\quad x\in(0,1),\\[2.0pt] w(t,0)=w(t,1)=0,\quad t\in[0,T],\end{array}\right$. (3.2)
where the inhomogeneous source term is given by
$\tilde{f}(t,x)=$ $f(t,x)-\partial_{t}\psi(t,x)+\partial_{x}(a(t,x)\partial_{x}\psi(t,x))-b(t,x)\partial_{x}\psi(t,x)-q(x)\psi(t,x)$
$=$ $f(t,x)-(1-x)g_{0}^{\prime}(t)-xg_{1}^{\prime}(t)+(\partial_{x}a(t,x)-b(t,x))(g_{1}(t)-g_{0}(t))$
$-q(x)\big((1-x)g_{0}(t)+xg_{1}(t)\big).$
The right-hand side $\tilde{f}\in L^{2}([0,T];L^{2}(0,1))$ under the assumptions $f\in L^{2}$ and $g_{0},g_{1}\in H^{1}([0,T])$, and the fact that the coefficients are bounded. Moreover, $\|\tilde{f}\|_{L^{2}}$ can be bounded by
$\|\tilde{f}\|_{L^{2}}\leq\|f\|_{L^{2}}+C\left(\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
for some constant $C>0$ depending only on $\|a\|_{W^{1,\infty}},\|b\|_{L^{\infty}},\|q\|_{L^{\infty}}$.
We now state the energy estimate.
Theorem 3.1.
Let $u$ solve the parabolic problem (1)–(1.3), with initial data $u_{0}\in L^{2}(0,1)$, source term $f\in L^{2}([0,T];L^{2}(0,1))$, and boundary data $g_{0},g_{1}\in H^{1}([0,T])$. Assume the coefficients satisfy
$a(t,x)\geq\alpha>0,\quad a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$
Then the solution $u\in L^{2}([0,T];H^{1}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$ satisfies:
$\|u\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}$
[…]
$\qquad\times\Big(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\Big),$ (3.3)
where $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
Proof.
As described above, we write $u=w+\psi$, where $w$ satisfies the homogeneous problem (3.2). Applying the energy estimate for $w$ from Theorem 2.2, we obtain:
$\|w\|_{L^{\infty}L^{2}}+\sqrt{\alpha}\|\partial_{x}w\|_{L^{2}L^{2}}+\|\sqrt{q}w\|_{L^{2}L^{2}}\leq Ce^{T\|\partial_{x}b\|_{L^{\infty}}}\left(\|u_{0}-\psi(0)\|_{L^{2}}+\|\tilde{f}\|_{L^{2}}\right).$ (3.4)
Since $\psi(0,x)=(1-x)g_{0}(0)+xg_{1}(0)$, we estimate:
$\|\psi(0)\|_{L^{2}}\leq C\left(|g_{0}(0)|+|g_{1}(0)|\right)\leq C\left(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}\right),$
hence,
$\|u_{0}-\psi(0)\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}+C(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
Next, since $u=w+\psi$, we use the triangle inequality:
$\|u\|_{L^{\infty}L^{2}}$ $\leq\|w\|_{L^{\infty}L^{2}}+\|\psi\|_{L^{\infty}L^{2}},$
$\|\partial_{x}u\|_{L^{2}L^{2}}$ $\leq\|\partial_{x}w\|_{L^{2}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}},$
$\|\sqrt{q}u\|_{L^{2}L^{2}}$ $\leq\|\sqrt{q}w\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}.$
Note that $\psi(t,x)\in\text{span}\{1,x\}\cdot g_{j}(t)$, so
$\|\psi\|_{L^{\infty}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}\leq C(T)\left(1+\|a\|^{1/2}_{L^{\infty}}+\|q\|^{1/2}_{L^{\infty}}\right)(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
Combining this with (3.4) yields (3.1), which completes the proof of Theorem 3.1. ∎
Corollary 3.2.
Let $u$ solve the non-homogeneous Dirichlet problem (1)–(1.3) with data
$f\in L^{2}([0,T];L^{2}(0,1)),\quad u_{0}\in L^{2}(0,1),\quad g_{0},g_{1}\in H^{1}([0,T]),$
and assume the coefficients satisfy:
$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
Then:
$\partial_{t}u,\,\partial_{x}(a\partial_{x}u)\in L^{2}([0,T];L^{2}(0,1)),$
and the following estimate holds:
$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial^{2}_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}\Big\}$
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
where the constant $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
Proof.
We decompose the solution as $u=w+\psi$, where $w$ solves the corresponding homogeneous Dirichlet problem with modified source
$\tilde{f}=f-\mathcal{L}\psi,\qquad\psi(t,x)=(1-x)g_{0}(t)+xg_{1}(t)$
is the boundary lifting.
Then, since $\partial_{t}\psi\in L^{2}([0,T];L^{2}(0,1))$ and $\partial_{x}(a\partial_{x}\psi)\in L^{2}([0,T];L^{2}(0,1))$, and the estimate from Proposition 2.4 applies to $w$, we have:
$\|\partial_{t}u\|_{L^{2}}\leq\|\partial_{t}w\|_{L^{2}}+\|\partial_{t}\psi\|_{L^{2}},\quad\|\partial_{x}(a\partial_{x}u)\|_{L^{2}}\leq\|\partial_{x}(a\partial_{x}w)\|_{L^{2}}+\|\partial_{x}(a\partial_{x}\psi)\|_{L^{2}}.$
Now, note that $\psi\in H^{1}([0,T];H^{1}(0,1))$, and all terms involving $\psi$ can be bounded in terms of $\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}$. Combining with the bound for $w$ from Proposition 2.4 yields the desired result. ∎
4. Very weak solutions
In this section, we consider the parabolic equation on $0,T]\times(0,1)$, where the diffusion coefficient, drift term, potential, initial data, and boundary data are modelled by distributions (e.g., Dirac-type functions). To treat such low regularity settings, we employ the concept of very weak solutions, formulated via moderate nets.
Let us assume that the potential $q$, initial condition $u_{0}$, and boundary data $g_{0},\,g_{1}$ are distributions. The diffusion coefficient $a$, drift term $b$, and source term $f$ are also allowed to be distributional in space and time.
Definition 2 (Moderate nets).
Let $\varepsilon\in(0,1]$, and consider nets of smooth approximations $(u_{\varepsilon})$, $(q_{\varepsilon})$, etc.
  • (i)
    A net $(u_{\varepsilon})\subset L^{2}([0,T];L^{2}(0,1))$ is said to be $L^{2}$-moderate if there exist constants $C>0$, $N\in\mathbb{N}$ such that
    $\|u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N}.$
  • (ii)
    A net $(q_{\varepsilon})\subset L^{\infty}(0,1)$ is $L^{\infty}$-moderate if there exist $C>0$, $N\in\mathbb{N}$ such that
    $\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\leq C\varepsilon^{-N}.$
Definition 3 (Coefficient moderateness).
Let $(a_{\varepsilon}),(b_{\varepsilon})$ be nets of smooth functions on $0,T]\times[0,1]$.
  • (i)
    The net $(a_{\varepsilon})\subset W^{1,\infty}([0,T]\times(0,1))$ is $W^{1,\infty}$-moderate if
    $\|a_{\varepsilon}\|_{W^{1,\infty}([0,T]\times(0,1))}\leq C\varepsilon^{-N}.$
  • (ii)
    The net $(b_{\varepsilon})\subset L^{\infty}([0,T];W^{1,\infty}(0,1))$ is log-$L^{\infty}(W^{1,\infty})$-moderate if
    $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\leq C|\log\varepsilon|^{N},\quad\|b_{\varepsilon}\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\leq C\varepsilon^{-N}.$
Definition 4 (Data moderation).
Let $u_{0,\varepsilon},f_{\varepsilon}$ be smooth approximations of initial and source data.
  • (i)
    The net $(u_{0,\varepsilon})\subset H^{1}(0,1)$ is $H^{1}$-moderate if
    $\|u_{0,\varepsilon}\|_{H^{1}(0,1)}\leq C\varepsilon^{-N}.$
  • (ii)
    The net $(f_{\varepsilon})\subset L^{2}([0,T];L^{2}(0,1))$ is $L^{2}$-moderate if
    $\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N}.$
Definition 5 (Distributional boundary data).
Let $g_{0,\varepsilon},g_{1,\varepsilon}\in C^{\infty}([0,T])$ be approximations of boundary values at $x=0$ and $x=1$.
We say the net $(g_{j,\varepsilon})$ is time-distributional moderate if there exist constants $C>0$, $N\in\mathbb{N}$ such that
$\|g_{j,\varepsilon}\|_{H^{1}([0,T])}\leq C\varepsilon^{-N},\quad\text{for }j=0,1.$
Remark 4.1.
The above assumptions are natural in the context of distributional coefficients, especially in one spatial dimension. Regularisations of distributions often form moderate nets. In particular, by the structure theorem for distributions (see, e.g., [16]), we have the inclusion
$\mathcal{D}^{\prime}(0,1)\subset\{\text{$L^{\infty}(0,1)$-moderate families}\}.$ (4.1)
This means that even if a solution to a boundary value problem on $(0,1)$ does not exist in the classical distributional sense, it may still exist as a net of $L^{\infty}$-moderate functions. This framework allows for a consistent treatment of singular coefficients and data.
As an example, let $f\in L^{2}(0,1)$, where $f:(0,1)\to\mathbb{C}$. We define the extension of $f$ by zero to all of $\mathbb{R}$ as:
$\widetilde{f}(t,x):=\begin{cases}f(x),&(x)\in(0,1),\\ 0,&(x)\in\mathbb{R}\setminus(0,1).\end{cases}$
[…]
Let $\psi\in C^{\infty}(\mathbb{R})$ be a standard mollifier with $\int_{\mathbb{R}}\psi=1$, and define the mollified net:
$\widetilde{f}_{\varepsilon}:=\widetilde{f}*\psi_{\varepsilon},\quad\text{where }\psi_{\varepsilon}(x):=\frac{1}{\varepsilon}\psi\left(\frac{x}{\varepsilon}\right),\quad\varepsilon\in(0,1].$
Then:
  • Each $\widetilde{f}_{\varepsilon}\in C^{\infty}(\mathbb{R})$,
  • The net $(\widetilde{f}_{\varepsilon})$ is $L^{p}(\mathbb{R})$-moderate for any $1\leq p\leq\infty$,
  • On the original domain, $\widetilde{f}_{\varepsilon}\to f$ in $L^{p}(0,1)$ as $\varepsilon\to 0$,
  • Outside $(0,1)$, the mollified functions decay, i.e., $\|\widetilde{f}_{\varepsilon}\|_{L^{p}(\mathbb{R}\setminus(0,1))}\to 0$.
Hence, the total approximation error can be estimated as:
$\|\widetilde{f}_{\varepsilon}-\widetilde{f}\|_{L^{p}(\mathbb{R})}^{p}\approx\|\widetilde{f}_{\varepsilon}-f\|_{L^{p}(0,1)}^{p}+\|\widetilde{f}_{\varepsilon}\|_{L^{p}(\mathbb{R}\setminus(0,1))}^{p}\to 0\quad\text{as }\varepsilon\to 0.$
This illustrates how mollification via extension by zero produces a smooth $L^{p}$-moderate net approximating a non-smooth function on $(0,1)$.
Similarly, one can apply mollification via extension by zero in the time variable to approximate functions on $[0,T]$ by smooth $L^{p}$-moderate nets.
Now, we define the concept of a very weak solution to the parabolic equation involving the divergence form with a diffusion coefficient, a drift-term and a potential, along with initial data and non-homogeneous Dirichlet boundary conditions
$\left\{\begin{array}[]{l}\partial_{t}u(t,x)-\partial_{x}(a\partial_{x}u)(t,x)+b\partial_{x}u(t,x)+q(x)u(t,x)=f(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ u(0,x)=u_{0}(x),\quad x\in(0,1),\\ u(t,0)=g_{0}(t),\,\,u(t,1)=g_{1}(t),\quad t\in[0,T].\end{array}\right$. (4.2)
Definition 6.
Assume that $a\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, and $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. A net of functions $(u_{\varepsilon})_{\varepsilon>0}$ is said to be a very weak solution to the initial/boundary value problem (4.2), if there exist an $L^{\infty}$-moderate regularisation $q_{\varepsilon}$ of $q$, a $W^{1,\infty}$-moderate regularisation $a_{\varepsilon}$ of $a$, a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation $b_{\varepsilon}$ of $b$, an $H^{1}$-moderate regularisation $u_{0,\varepsilon}$ of $u_{0}$, time-distributional moderate regularisations $g_{0,\varepsilon}$ and $g_{1,\varepsilon}$ of $g_{0}$ and $g_{1}$ respectively, and an $L^{2}$-moderate regularisation $f_{\varepsilon}$ of $f$, such that the regularised problem
$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),\\ \hfill(t,x)\in[0,T]\times(0,1),\\ u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\quad x\in(0,1),\\ u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\quad t\in[0,T],\end{array}\right$. (4.3)
holds for all $\varepsilon\in(0,1]$, and the net $(u_{\varepsilon})$ is moderate in the sense that there exist constants $C>0$, $N\in\mathbb{N}$ such that
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},\quad\text{for all }\varepsilon\in(0,1].$
Theorem 4.1 (Existence of Very Weak Solutions).
Let the coefficients $q\in\mathcal{D}^{\prime}(0,1)$, $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, and the boundary data $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. Then, the initial-boundary value problem (4.2) admits a very weak solution in the sense of Definition 6.
Proof.
Since equation (4.2) involves products of distributions, we replace it with a regularised problem using smooth approximations of all data and coefficients.
Let $(q_{\varepsilon})$ be an $L^{\infty}$-moderate regularisation of $q$, $(a_{\varepsilon})$ a $W^{1,\infty}$-moderate regularisation of $a$, with $a_{\varepsilon}\geq\alpha>0$, $(b_{\varepsilon})$ a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation of $b$, $(u_{0,\varepsilon})$ an $H^{1}$-moderate regularisation of $u_{0}$, $(g_{0,\varepsilon}),(g_{1,\varepsilon})$ time-distributional moderate regularisations of $g_{0},g_{1}$, and $(f_{\varepsilon})$ an $L^{2}$-moderate regularisation of $f$. Then, for each $\varepsilon\in(0,1]$, consider the regularised problem (4.3).
For each fixed $\varepsilon\in(0,1]$, the coefficients and data in (4.3) are smooth, and the diffusion coefficient $a_{\varepsilon}$ is uniformly elliptic. Hence, by classical parabolic theory and the energy estimate (see Corollary 3.2), there exists a unique weak solution
$u_{\varepsilon}\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1)).$
Moreover, Corollary 3.2 implies the energy estimate
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\left(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\right).$ (4.4)
Using the moderateness assumptions for all regularised nets, the right-hand side is bounded by $C\varepsilon^{-N}$ for some $N\in\mathbb{N}$, uniformly in $\varepsilon\in(0,1]$. Thus,
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},$
so $(u_{\varepsilon})$ is moderate in $H^{1}L^{2}\cap L^{2}H^{2}$, and therefore, by Definition 6, $(u_{\varepsilon})$ is a very weak solution to (4.2). ∎
Definition 7 (Negligibility).
Let $(u_{\varepsilon})$, $(\tilde{u}_{\varepsilon})$ be two nets in $L^{\infty}([0,T];L^{2}(0,1))$. We say that the net $(u_{\varepsilon}-\tilde{u}_{\varepsilon})$ is $L^{2}$-negligible if, for every $N\in\mathbb{N}$ there exist a constant $C>0$ such that
$\|u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)\|_{L^{2}}\leq C\varepsilon^{N},$
for all $\varepsilon\in(0,1]$ uniformly in $t\in[0,T]$. The constant $C$ may depends on $N$ but not on $\varepsilon$.
Let us consider the $\varepsilon$-parameterised problems”:
$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)\\ \hfill=f_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\quad t\in[0,T],\end{array}\right$. (4.5)
and
$\left\{\begin{array}[]{l}\partial_{t}\tilde{u}_{\varepsilon}(t,x)-\partial_{x}(\tilde{a}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon})(t,x)+\tilde{b}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon}(t,x)+\tilde{q}_{\varepsilon}(x)\tilde{u}_{\varepsilon}(t,x)\\ \hfill=\tilde{f}_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ \tilde{u}_{\varepsilon}(0,x)=\tilde{u}_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ \tilde{u}_{\varepsilon}(t,0)=\tilde{g}_{0,\varepsilon}(t),\quad\tilde{u}_{\varepsilon}(t,1)=\tilde{g}_{1,\varepsilon}(t),\quad t\in[0,T].\end{array}\right$. (4.6)
Definition 8 (Uniqueness of the very weak solution).
Let $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, $g_{0},\,g_{1}\in\mathcal{D}^{\prime}([0,T])$. We say that the initial/boundary problem (4.2) has a unique very weak solution, if for all $W^{1,\infty}(L^{2})$, log-$W^{1,\infty}$-moderate nets $a_{\varepsilon}$, $\tilde{a}_{\varepsilon}$, $b_{\varepsilon}$, $\tilde{b_{\varepsilon}}$, such that $(a_{\varepsilon}-\tilde{a}_{\varepsilon})$, $(b_{\varepsilon}-\tilde{b}_{\varepsilon})$ are $W^{1,\infty}$-negligible and log-$L^{\infty}(W^{1,\infty})$-negligible; $L^{\infty}$-moderate nets $q_{\varepsilon}$, $\tilde{q}_{\varepsilon}$, such that $(q_{\varepsilon}-\tilde{q}_{\varepsilon})$ are $L^{\infty}$-negligible; time-distributional moderate nets $g_{0,\varepsilon},\,g_{1,\varepsilon},\,\tilde{g}_{0,\varepsilon},\,\tilde{g}_{1,\varepsilon}$, such that $(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})$, $(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})$ are $H^{1}$-negligible; for all $L^{2}$-moderate nets $f_{\varepsilon},\,\tilde{f}_{\varepsilon}$, such that $(f_{\varepsilon}-\tilde{f}_{\varepsilon})$ are $L^{2}$-negligible; for all $H^{1}$-moderate nets $u_{0,\varepsilon},\,\tilde{u}_{0,\varepsilon}$ such that $(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})$ is $H^{1}$-negligible, we have that $u_{\varepsilon}-\tilde{u}_{\varepsilon}$ is $H^{1}(L^{2})\cap L^{2}(H^{2})$-negligible.
Theorem 4.2 (Uniqueness of the very weak solution).
Let the coefficient $q$ be distribution on $(0,1)$, the diffusion $a$, drift term $b$ be distributions on $0,T]\times(0,1)$, and the boundary data $g_{0},\,g_{1}\in\mathcal{D}^{\prime}([0,T])$. Then the very weak solution to the initial/boundary problem (4.2) is unique.
Proof.
Let us denote by $u_{\varepsilon}$ and $\tilde{u}_{\varepsilon}$ the families of solutions for the initial/boundary problems (4.5) and (4.6) correspondingly. Let $U_{\varepsilon}$ represents the difference between these nets $U_{\varepsilon}(t,\cdot):=u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)$ then $U_{\varepsilon}$ solves the equation
$\left\{\begin{array}[]{l}\partial_{t}U_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}U_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}U_{\varepsilon}(t,x)+q_{\varepsilon}(x)U_{\varepsilon}(t,x)\\ \hfill=F_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ U_{\varepsilon}(0,x)=(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})(x),\,\,\,x\in(0,1),\\ U_{\varepsilon}(t,0)=(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})(t),\,U_{\varepsilon}(t,1)=(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})(t),\,\,t\in[0,T],\end{array}\right$. (4.7)
where we set
$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\left(\big(\tilde{a}_{\varepsilon}(t,x)-a_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}\right)(t,x)+\big(\tilde{b}_{\varepsilon}(t,x)-b_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}(t,x)$
$+\big(\tilde{q}_{\varepsilon}(x)-q_{\varepsilon}(x)\big)\tilde{u}_{\varepsilon}(t,x)+f_{\varepsilon}(t,x)-\tilde{f}_{\varepsilon}(t,x)$
for the source term to the initial/boundary problem (4.7).
Taking the $H^{1}(L^{2})\cap L^{2}(H^{2})$-norm of the $U_{\varepsilon}$ by using (4), we obtain
$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\quad\times\left(\|U_{\varepsilon}(0,\cdot)\|_{L^{2}}+\|F_{\varepsilon}\|_{L^{2}}+\|U_{\varepsilon}(\cdot,0)\|_{H^{1}}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}}\right)\Bigg).$
For the source term $F_{\varepsilon}$, we have
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\partial_{x}((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\tilde{b}_{\varepsilon}-b_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\tilde{q}_{\varepsilon}-q_{\varepsilon})\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$ (4.8)
For the first term in the right-hand side we have
$\partial_{x}\big((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\big)=\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}+(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}^{2}\tilde{u}_{\varepsilon}.$
Hence
$\|\partial_{x}\cdot((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|$ $_{L^{2}([0,T];L^{2}(0,1))}\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
By Proposition 2.4 and Corollary 3.2, this quantity is finite and moderate. Then, we have
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
Combining the initial condition of (4.7), we obtain
$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\Big(\|u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
Considering the negligibility of the nets $\tilde{a}_{\varepsilon}-a_{\varepsilon}$, $\tilde{b}_{\varepsilon}-b_{\varepsilon}$, $q_{\varepsilon}-\tilde{q}_{\varepsilon}$, $u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}$, $g_{\varepsilon}-\tilde{g}_{\varepsilon}$, $f_{\varepsilon}-\tilde{f}_{\varepsilon}$, we get
$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\qquad\leq\frac{C_{0}}{\alpha}\Bigg(\varepsilon^{-N_{0}}\Big(C_{1}\varepsilon^{-N_{1}}+C_{2}\varepsilon^{-N_{2}}+\frac{1}{\sqrt{\alpha}}\left(C_{3}\varepsilon^{-N_{3}}+C_{4}\varepsilon^{-N_{4}}+C_{5}\varepsilon^{-N_{5}}\right)\Big)$
$\qquad\times\Big(1+\frac{1}{\sqrt{\alpha}}\left(C_{6}\varepsilon^{-N_{6}}+C_{7}\varepsilon^{-N_{7}}\right)+C_{8}\varepsilon^{-N_{8}}\Big)\big(C_{9}\varepsilon^{N_{9}}+C_{10}\varepsilon^{N_{10}}\varepsilon^{-N_{11}}$
$\qquad+C_{11}\varepsilon^{N_{12}}\varepsilon^{-N_{13}}+C_{12}\varepsilon^{N_{14}}\varepsilon^{-N_{15}}+C_{13}\varepsilon^{N_{16}}\varepsilon^{-N_{17}}+C_{14}\varepsilon^{N_{18}}+C_{15}\varepsilon^{N_{19}}+C_{16}\varepsilon^{N_{20}}\big)\Bigg)$
for some $C_{j}>0$, j=0,,16j=0,...,16, $N_{k}\in\mathbb{N}$, $k=0,\ldots,8,\,11,\,13,\,15,\,17$, and all
$N_{9},\,N_{10},\,N_{12},\,N_{14},\,N_{16},\,N_{18},\,N_{19},\,N_{20}\in\mathbb{N}$. Then, for some $C_{D}>0$ and all $K\in\mathbb{N}$
$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\leq C_{D}\varepsilon^{K},$
and this completes the proof of Theorem 4.2. ∎
Theorem 4.3 (Consistency).
Assume that $a\in W^{1,\infty}([0,T];W^{1,\infty}(0,1))$,
$b\in L^{\infty}([0,T];W^{1,\infty}(0,1))$, $q\in L^{\infty}(0,1)$, $g_{0},\,g_{1}\in H^{1}([0,T])$, and let $a_{\varepsilon}$ be any $W^{1,\infty}$-regularisation of $a$, $b_{\varepsilon}$ be any log-$L^{\infty}(W^{1,\infty})$-regularisation of $b$, $(q_{\varepsilon})$ be any $L^{\infty}$-regularisation of $q$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$ are any time-distributional regularisations of $g_{0}$ and $g_{1}$ respectively, that is
$\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
$\|q_{\varepsilon}-q\|_{L^{\infty}(0,1)}\to 0,\quad\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0$
as $\varepsilon\to 0$. Let the initial condition satisfies $u_{0}\in H^{1}(0,1)$. Let $u$ be a very weak solution of the initial/boundary problem (4.2). Then for any families $a_{\varepsilon}$, $b_{\varepsilon}$, $q_{\varepsilon}$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$, $u_{0,\varepsilon}$ and $f_{\varepsilon}$ such that
$\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
$\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\to 0,\quad\|u_{0}-u_{0,\varepsilon}\|_{H^{1}(0,1)}\to 0,$
$\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
as $\varepsilon\to 0$, any representative $(u_{\varepsilon})$ of the very weak solution converges as
$\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{H^{1}([0,T];L^{2}(0,1))}+\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
for $\varepsilon\to 0$ to the unique classical solution $u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1))$ of the initial/boundary problem (4.2) given by Theorem 3.1.
Proof.
Let us denote $\tilde{U}_{\varepsilon}$ as the difference between $u$ and $u_{\varepsilon}$, that is
$\tilde{U}_{\varepsilon}(t,x):=u(t,x)-u_{\varepsilon}(t,x).$
Then the net $\tilde{U}_{\varepsilon}$ is a solution to the initial/boundary problem
$\left\{\begin{array}[]{l}\partial_{t}\tilde{U}_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon})(t,x)+b_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon}+q_{\varepsilon}(x)\tilde{U}_{\varepsilon}(t,x)=F_{\varepsilon}(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ \tilde{U}_{\varepsilon}(0,x)=(u_{0}-u_{0,\varepsilon})(x),\quad x\in(0,1),\\ \tilde{U}_{\varepsilon}(t,0)=(g_{0}-g_{0,\varepsilon})(t),\quad\tilde{U}_{\varepsilon}(t,1)=(g_{1}-g_{1,\varepsilon})(t),\quad t\in[0,T],\end{array}\right$. (4.9)
where
$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\cdot\left(\big(a_{\varepsilon}(t,x)-a(t,x)\big)\partial_{x}u\right)(t,x)+\big(b_{\varepsilon}(t,x)-b(t,x)\big)\cdot\partial_{x}u(t,x)$
$+\big(q_{\varepsilon}(x)-q(x)\big)u(t,x)+f(t,x)-f_{\varepsilon}(t,x)$
Analogously to Theorem 4.2 we have that
$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\Big(\|u_{0}-{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|{a}_{\varepsilon}-a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}({a}_{\varepsilon}-a)\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|{b}_{\varepsilon}-b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|{q}_{\varepsilon}-q\|_{L^{\infty}(0,1)}\|{u}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f-{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
Since
$\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
$\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\to 0,\quad\|u_{0}-u_{0,\varepsilon}\|_{H^{1}(0,1)}\to 0,$
$\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])},\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])},\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
for $\varepsilon\to 0$ and $u$ is a very weak solution of the initial/boundary problem (4.2), we get
$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
for $\varepsilon\to 0$. This proves Theorem 4.3.
5. Numerical Experiments
In this Section, we carry out numerical experiments of the parabolic equations with singular coefficients in the one-dimensional case. More precisely, we consider the initial–boundary value problem
$\partial_{t}u(t,x)-\partial_{x}\left(a(t,x)\partial_{x}u(t,x)\right)+b(t,x)\partial_{x}u(t,x)+q(x)u(t,x)=f(t,x),$
$\quad(t,x)\in[0,T]\times(0,1),$ (5.1)
with initial condition
$u(0,x)=u_{0}(x),\quad x\in(0,1),$ (5.2)
and inhomogeneous Dirichlet boundary conditions
$u(t,0)=g_{0}(t),\quad u(t,1)=g_{1}(t),\quad t\in[0,T].$ (5.3)
In this work, we are interested in the singular cases of the coefficients $a(t_{0},x)$, $b(t,x_{0})$, $q(x),\,u_{0}(x),\,g_{0}(t),\,g_{1}(t)$. Even, we can allow them to be distributional, in particular, to have $\delta$-like singularities. As it was theoretically outlined in [22] and [23], we start to analyse our problem by regularising a distributional valued functions $a(t,x),b(t,x),q(t),u_{0}(x),g_{0}(t),g_{1}(t)$ by a parameter $\varepsilon$, that is, we set
$a_{\varepsilon}(x):=(a*\varphi_{\varepsilon})(x),\quad b_{\varepsilon}(t):=(b*\varphi_{\varepsilon})(t),\quad q_{\varepsilon}(x):=(q*\varphi_{\varepsilon})(x),\quad$ (5.4)
$g_{0,\varepsilon}(t):=(g_{0}*\varphi_{\varepsilon})(t),\quad g_{1,\varepsilon}(t):=(g_{1}*\varphi_{\varepsilon})(t)$ (5.5)
as the convolution with the mollifier
$\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(x/\varepsilon),$ (5.6)
where $\varphi(x)=c\exp{\left(\frac{1}{x^{2}-1}\right)}$ for $|x|<1,$ and $\varphi(x)=0$ otherwise. Here $c\simeq 2.2523$ to get $\int\limits_{-\infty}^{\infty}\varphi(x)dx=1.$ For $t$ we also use the same mollifier, which is dependent on $t$. Then, instead of (1) we consider the regularised problem
$\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}\left(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)\right)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),$
$\quad(t,x)\in[0,T]\times(0,1),$ (5.7)
with the Cauchy data
$u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),$
for all $x\in\left[0,1\right].$ Here, we put
$u_{0}(x)=exp\left(\frac{1}{(x-0.5)^{2}+0.025}\right)$
For the coefficients $a,b,q,g_{0},g_{1}$ we consider the following cases, with $\delta$ denoting the standard Dirac’s delta-distribution: First, we study the following model situation:
  • Case 1, $a=1,\quad b=1,\quad q=1,\quad g_{0}=g_{1}=0$
  • Case 2, $a=1+\delta(x-0.45),\quad b=1,\quad q=1,\quad g_{0}=g_{1}=0$
  • Case 3, $a=1,\quad b=1+\delta(t-0.5),\quad q=1,\quad g_{0}=g_{1}=0$
  • Case 4, $a=1,\quad b=1,\quad q=1+\delta(x-0.45),\quad g_{0}=g_{1}=0$
  • Case 5, $a=1,\quad b=1,\quad q=1,\quad g_{0}=0\quad g_{1}=\delta(t-0.45),$
For comparison with other cases in figure 1, we illustrate the regular case when all the coefficients are constant, such that $a=1.0,b=1.0,q=1,g_{0}=g_{1}=0$, the behaviour of the heat function $u(t,x)$ at different time steps.
Figure 1. Case 1
The evolution of the heat distribution in Figure 1 displays smooth temporal decay accompanied by a rightward shift, which is a direct consequence of the transport effected by the $u_{x}$ term.
Figure 2 illustrates the behaviour of the heat function $u(t,x)$, which serves as the solution to equation 5, analysed for $\varepsilon$=0.3, 0.1, 0.05, 0.031 and 0.003 in Case 2. In this scenario, the coefficient $a(t,x)$ is represented by a Dirac delta function supported at the point $x=0.45$.
Refer to captionRefer to caption
Figure 2. Case 2
Figure 2 illustrates the temporal decay of the heat distribution $u(t,x)$. A pronounced singularity is observed at the location of the Dirac delta source ($x=0.45$). The magnitude of this singularity is controlled by the parameter $\varepsilon$, with smaller values of $\varepsilon$ leading to a more intense and localised peak.
Figure 3 illustrates the behaviour of the heat function $u(t,x)$, which serves as the solution to equation 5, analysed for $\varepsilon$=0.001 in Case 3. In this scenario, the coefficient $b(t,x)$ is represented by a Dirac delta function supported at point t=0.5. On the right side analyzed for various $\varepsilon$=0.1 ,0.01, 0.005, 0.002 and 0.0001 in Case 3.
Refer to captionRefer to caption
Figure 3. Case 3
Figure 3 shows the evolution of the heat distribution, which decays smoothly until the singularity time $t=0.25$. At this instant, the solution exhibits a discontinuous jump to the right, a phenomenon driven by the advective term $b(t,x)u_{x}(t,x)$. Following the jump, the distribution resumes its smooth decay.
Furthermore, the right figure demonstrates that the magnitude of this jump is inversely proportional to the parameter $\varepsilon$; smaller values of $\varepsilon$ result in a larger jump.
Figure 4 analyses the heat distribution under a Dirac delta potential $q(x)$ at $x=0.6$. On the left, the evolution of the distribution is shown for a fixed $\varepsilon=0.005$. On the right, the results at the final time are compared for different values of $\varepsilon$.
Refer to captionRefer to caption
Figure 4. Case 4
Figure 4 shows that the overall heat distribution decays over time. Concurrently, the Dirac delta potential at $x=0.6$ induces a localised cooling effect. As illustrated on the left, this cooling effect occurs more rapidly for smaller values of $\varepsilon$.
Figure 5 illustrates the heat distribution subject to an inhomogeneous Dirichlet boundary condition with $g_{1}(t)$ modelled by a Dirac delta function at $t=0.25$. The left figure shows the distribution’s evolution for a fixed $\varepsilon=0.005$. The right figure compares the profiles at the final time for various values of $\varepsilon$.
Refer to captionRefer to caption
Figure 5. Case 5
As in previous cases, the heat distribution in Figure 5 decays over time. At $t=0.25$, the Dirac delta boundary condition activates, causing a localised heating effect near the right boundary. This temperature increase is transient; after several time steps, the boundary temperature decreases. The right panel shows that this localised heating occurs more rapidly for smaller values of $\varepsilon$.

### v2

@@
 Abstract.
-We investigate linear parabolic equations in divergence form with singular coefficients and non-smooth boundary data.
-When the diffusion, drift, or potential terms, as well as the initial or boundary conditions, are distributions rather than functions, classical and weak solution concepts become inadequate due to the ill-posedness of products involving distributions.
-To overcome this, we introduce a framework of very weak solutions based on regularization techniques and the theory of moderate nets.
-Existence of very weak solutions is established under minimal regularity assumptions.
-We further prove consistency with classical solutions when the data are smooth and demonstrate uniqueness via negligibility arguments.
-Finally, we present numerical computations that illustrate the robustness of the very weak solution framework in handling highly singular inputs, including delta-type potentials and distributional boundary traces.
+We investigate linear parabolic equations in divergence form with singular coefficients and non-smooth initial–boundary data.
+When the diffusion, drift, or potential terms, as well as the source term and boundary conditions, are distributions rather than functions, classical and weak solution concepts become inadequate, since products involving distributions are not well defined in general.
+To address this difficulty, we introduce a framework of very weak solutions based on regularisation procedures and the theory of moderate nets.
+Under minimal regularity assumptions, we establish existence of very weak solutions and prove uniqueness via negligibility arguments.
+Moreover, in the smooth-data regime, we show consistency with classical weak solutions.
+Finally, we present numerical experiments illustrating the effectiveness of the proposed framework for highly singular inputs, including delta-type potentials and distributional boundary traces.
 Key words and phrases:
-parabolic equation, divergence form, shift term, energy methods, Galerkin approximation, singular coefficients, regularisation, very weak solution.
+parabolic equation, divergence form, drift term, energy methods, Galerkin approximation, singular coefficients, regularisation, very weak solution.
 2020 Mathematics Subject Classification:
@@
 $u(t,0)=g_{0}(t),\quad u(t,1)=g_{1}(t),\quad t\in[0,T].$ (1.3)
-Here, $a(t,x)$ is the (possibly singular) diffusion coefficient associated with the divergence term, $b(t,x)$ represents the drift field, $q(x)$ is a spatial potential, $f(t,x)$ is an external source term, and $u_{0}(x)$, $g_{0}(t)$, and $g_{1}(t)$ are the prescribed initial and boundary data, respectively.
+Here $a(t,x)$ denotes the diffusion coefficient, $b(t,x)$ the drift field, $q(x)$ a potential term, and $f(t,x)$ an external source.
+The functions $u_{0}$, $g_{0}$, and $g_{1}$ prescribe the initial and boundary data.
 The well-posedness of such parabolic problems with regular coefficients and boundary data is a classical subject in the theory of partial differential equations (see, for instance, [9, 10, 11, 16]).
-However, when the diffusion coefficient $a$, the drift term $b$, the potential $q$, the source term $f$, the initial data $u_{0}$, or the boundary data $g_{0},g_{1}$ are allowed to be non-regular (e.g., involving Dirac delta functions or other distributions), the standard functional frameworks break down.
+However, when the diffusion coefficient $a$, the drift term $b$, the potential $q$, the source term $f$, the initial data $u_{0}$, or the boundary data $g_{0},g_{1}$ are allowed to be distributional (e.g., involving Dirac delta functions or other distributions), the standard functional frameworks break down.
 In such cases, the product of distributions becomes undefined, and well-posedness in the classical or even weak sense fails to hold.
-From an analytical standpoint, the study of PDEs with distributional coefficients lies at the frontier of modern functional analysis and the theory of generalized functions.
-The classical theory, built upon Sobolev spaces and variational methods, becomes inadequate in the presence of distributional multiplicative terms due to Schwartz’s impossibility theorem [28].
+From an analytical standpoint, PDEs with distributional coefficients pose substantial challenges within the theory of generalized functions.
+Classical approaches based on Sobolev spaces and variational methods become inadequate, since products of distributions are not well defined in general, as shown by Schwartz’s impossibility theorem [32].
 Similar challenges arise in hyperbolic equations.
 To address them, the notion of very weak solutions was introduced in [22], enabling analysis of second-order equations with singular coefficients.
-This approach has since been developed in numerous works addressing various types of singularities and equations (see [1, 2, 3, 4, 5, 9, 10, 11, 12, 17, 18, 20, 22, 23, 29]).
+This approach has since been developed in numerous works addressing various types of singularities and equations (see [1, 2, 3, 4, 5, 6, 7, 8, 12, 13, 14, 15, 20, 21, 24, 26, 27, 33]).
 Recent works on elliptic, hyperbolic and parabolic equations for the Sturm-Liouville operator with singular coefficients [25, 28, 29, 30] have successfully employed separation of variables to construct very weak solutions.
-In particular, the parabolic case presents additional challenges due to the time evolution aspect, which complicates the treatment of singularities in a dynamic context.
-The extension of the very weak solution framework to such settings is both analytically rich and methodologically nontrivial.
-In our approach, we employ regularisation techniques to treat singular coefficients and data, and we derive uniform a priori estimates to control the behaviour of solution sequences.
-Moreover, we establish consistency with classical solutions in the smooth-data regime, ensuring the robustness and completeness of our methodology.
+The parabolic case presents additional difficulties, since singularities interact with the time evolution and uniform estimates must be established already at the level of the regularised problems.
+Extending the very weak solution framework to this setting therefore requires additional analytical tools.
+In our approach, we regularise the coefficients and data and derive uniform a priori bounds to control the resulting solution families.
+Moreover, we prove consistency with classical weak solutions in the smooth-data regime, which confirms the validity of the proposed methodology.
 Singularities in model coefficients and data arise naturally in many physical and engineering applications.
@@
 Our results provide justification for numerical approximations by showing that very weak solutions are limits of regularized problems for which standard numerical schemes (e.g., finite differences or finite elements) are applicable.
-The purpose of this work is to extend the framework of very weak solutions to the linear parabolic equation (1)–(1.3) in the presence of both spatial and time-dependent singularities.
-Our analysis accommodates singularities in the diffusion coefficient $a$, drift term $b$, potential $q$, source term $f$, initial data $u_{0}$, and boundary data $g_{0},g_{1}$, by regularising the problematic terms and constructing solution sequences with uniform estimates.
-This strategy allows us to prove the existence, uniqueness, and consistency of very weak solutions with classical solutions when the data are smooth.
-In a recent work by the authors [27], the analysis of the Klein–Gordon equation with singular coefficients was carried out under inhomogeneous Dirichlet boundary conditions.
-The analytical techniques developed in that study are applicable to the present problem, with appropriate modifications.
-In this work, we extend the analytical framework to parabolic partial differential equations with singular coefficients and inhomogeneous Dirichlet boundary conditions in the one-dimensional spatial setting.
-Our formulation accommodates singularities in all components of the model—including diffusion, drift, potential, source terms, and boundary data—through a regularization strategy.
-We establish well-posedness in the very weak sense under this framework.
-Additionally, we present numerical simulations that illustrate the behaviour of the solutions.
-Our methodology builds upon and adapts analytical techniques previously developed for the Klein–Gordon equation with singularities [27].
-The primary novelty of this work lies in the analysis of a parabolic equation with singular coefficients and a source term, under inhomogeneous Dirichlet boundary conditions — a setting which, to the best of our knowledge, has not yet been investigated in the literature.
-This study provides a robust theoretical framework for future analytical and numerical developments in this direction.
+The aim of this work is to extend the very weak solution framework to the parabolic problem (1)–(1.3) in the presence of singular coefficients and inhomogeneous Dirichlet boundary data.
+We allow distributional diffusion and drift coefficients, singular potentials and source terms, as well as non-smooth initial and boundary conditions.
+By regularising the coefficients and data and establishing uniform a priori estimates, we prove existence and uniqueness of very weak solutions, together with consistency with classical weak solutions in the smooth regime.
+Finally, we provide numerical simulations illustrating the behaviour of regularised solutions under delta-type singularities.
+The present work continues our recent analysis of the Klein–Gordon equation with singular coefficients under inhomogeneous Dirichlet boundary conditions [31].
+Many of the analytical techniques developed there, including the regularisation strategy and the derivation of uniform a priori estimates, provide the foundation for the current study.
+In contrast to the hyperbolic setting, the parabolic case introduces additional difficulties, since singularities interact with the time evolution and require new arguments.
+To the best of our knowledge, well-posedness for parabolic divergence-form equations with distributional coefficients and non-homogeneous Dirichlet boundary data has not previously been established within the very weak solution framework.
+Thus, the primary novelty of this work lies in the analysis of a parabolic equation with singular diffusion, drift, and potential terms, together with a singular source term, under inhomogeneous Dirichlet boundary conditions.
+The paper is organised as follows.
+In Section 2 we recall the energy estimates for the homogeneous Dirichlet problem.
+Section 3 contains the existence, uniqueness, and consistency results for very weak solutions under inhomogeneous Dirichlet boundary conditions.
+Section 4 is devoted to numerical experiments.
 2. Homogeneous Dirichlet boundary condition
@@
 where the diffusion and drift coefficients, respectively $a,b\in L^{\infty}([0,T];L^{\infty}(0,1))$ with $a(t,x)\geq\alpha>0$, the potential $q\in L^{\infty}(0,1)$, the source term $f\in L^{2}\big([0,T];L^{2}(0,1)\big)$, and the initial data $u_{0}\in L^{2}(0,1)$.
+Since $q$ may change sign and coercivity may fail, we work in the complex Hilbert spaces $L^{2}(0,1;\mathbb{C})$ and $H_{0}^{1}(0,1;\mathbb{C})$ endowed with the standard sesquilinear $L^{2}$ inner product
+$(u,v)_{L^{2}(0,1)}:=\int_{0}^{1}u(x)\,\overline{v(x)}\,dx,\qquad\|u\|_{L^{2}(0,1)}^{2}:=(u,u)_{L^{2}(0,1)}.$
+All weak formulations below are understood with this convention.
 To build a weak solution of the problem (2.1) we use Galerkin’s method, as in [31].
@@
 Definition 1 (Weak solution).
-A function $u\in L^{2}(0,T;H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ is said to be a weak solution to the problem (2.1), if the following identity holds for all test functions $\varphi\in H^{1}_{0}(0,1)$ and almost every $t\in[0,T]$:
-$\frac{d}{dt}\int_{0}^{1}u(t,x)\varphi(x)\,dx$ $+\int_{0}^{1}a(t,x)\,\partial_{x}u(t,x)\,\partial_{x}\varphi(x)\,dx+\int_{0}^{1}b(t,x)\,\partial_{x}u(t,x)\,\varphi(x)\,dx$
-$+\int_{0}^{1}q(x)\,u(t,x)\,\varphi(x)\,dx=\int_{0}^{1}f(t,x)\,\varphi(x)\,dx.$ (2.2)
+A function
+$u\in L^{2}(0,T;H^{1}_{0}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),\qquad\partial_{t}u\in L^{2}(0,T;H^{-1}(0,1)),$
+is called a weak solution of (2.1) if $u(0)=u_{0}$ in $L^{2}(0,1)$ and for all $\varphi\in H^{1}_{0}(0,1)$ and for almost every $t\in(0,T)$,
+$\langle\partial_{t}u(t),\varphi\rangle_{H^{-1},H^{1}_{0}}$ $+\int_{0}^{1}a(t,x)\,\partial_{x}u(t,x)\,\overline{\partial_{x}\varphi(x)}\,dx+\int_{0}^{1}b(t,x)\,\partial_{x}u(t,x)\,\overline{\varphi(x)}\,dx$
+$+\int_{0}^{1}q(x)\,u(t,x)\,\overline{\varphi(x)}\,dx=\int_{0}^{1}f(t,x)\,\overline{\varphi(x)}\,dx.$ (2.2)
 2.1. Galerkin Approximation
@@
 This is a linear system of ODEs for the vector $\mathbf{d}_{m}(t)=(d_{m}^{1}(t),\ldots,d_{m}^{m}(t))^{\top}$.
-Since all coefficients are measurable and bounded on $[0,T]$, standard ODE theory ensures the existence and uniqueness of a solution $\mathbf{d}_{m}\in C^{1}([0,T];\mathbb{R}^{m})$, from which the function $u_{m}$ follows via (2.4).
+Since all coefficients are measurable and bounded on $[0,T]$, standard ODE theory ensures the existence and uniqueness of a solution $\mathbf{d}_{m}\in C^{1}([0,T];\mathbb{C}^{m})$, from which the function $u_{m}$ follows via (2.4).
 ∎
@@
 Let $u_{m}$ be the Galerkin approximation defined in (2.4), and assume that the coefficients satisfy
-$a(t,x)\geq\alpha>0,\quad a,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$ (2.8)
+$a(t,x)\geq\alpha>0,\quad a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1),$ (2.8)
+and
+$q=q_{+}-q_{-},\quad q_{+}\geq 0,\quad q_{-}:=\max\{-q,0\}.$
 Then, for all $m\in\mathbb{N}$, the following energy estimate holds:
-$\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|\partial_{t}u_{m}\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
+$\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+}}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|\partial_{t}u_{m}\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\cdot\exp{\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}}$
+$\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}([0,T]\times(0,1))}+\|b\|_{L^{\infty}([0,T]\times(0,1))}\right)+\|q\|_{L^{\infty}(0,1)}\right)$
+$\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\right),$ (2.9)
+where $C>0$ depends only on $T$ and the spatial domain $(0,1)$, and is independent of $m$.
+All dependence on the coefficients $a,b,q$ (and on the data $u_{0},f$) is explicitly displayed on the right-hand side of (2.2).
+Proof.
+We test the weak formulation (2.6) with $u_{m}$. Taking real parts, we obtain
+$\operatorname{Re}(\partial_{t}u_{m},u_{m})+\operatorname{Re}(a\,\partial_{x}u_{m},\partial_{x}u_{m})+\operatorname{Re}(b\,\partial_{x}u_{m},u_{m})+\operatorname{Re}(qu_{m},u_{m})=\operatorname{Re}(f,u_{m}).$
+The time derivative term satisfies
+$\operatorname{Re}(\partial_{t}u_{m},u_{m})=\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
+Using the ellipticity condition $a\geq\alpha>0$, the diffusion term yields
+$\operatorname{Re}(a\,\partial_{x}u_{m},\partial_{x}u_{m})\geq\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}.$
+For the drift term, integration by parts (using the homogeneous Dirichlet boundary condition) gives
+$\operatorname{Re}(b\,\partial_{x}u_{m},u_{m})=-\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}).$
+The potential term can be written as
+$\operatorname{Re}(qu_{m},u_{m})=(q,|u_{m}|^{2}).$
+We decompose the potential as $q=q_{+}-q_{-}$, where $q_{+}\geq 0$ and $q_{-}:=\max\{-q,0\}$. Hence,
+$(q,|u_{m}|^{2})=(q_{+},|u_{m}|^{2})-(q_{-},|u_{m}|^{2}).$
+Combining the above identities, we arrive at the energy inequality
+$\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}$ $+\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+(q_{+},|u_{m}|^{2})$
+$\leq\operatorname{Re}(f,u_{m})+(q_{-},|u_{m}|^{2})+\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}).$
+We estimate the individual terms as follows. First,
+$(q_{+},|u_{m}|^{2})=\int_{0}^{1}q_{+}|u_{m}|^{2}\,dx=\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}(0,1)}.$
+Similarly,
+$(q_{-},|u_{m}|^{2})=\|\sqrt{q_{-}}u_{m}\|^{2}_{L^{2}(0,1)}\leq\|q_{-}\|_{L^{\infty}(0,1)}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
+For the drift contribution, we have
+$\frac{1}{2}(\partial_{x}b,|u_{m}|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
+Finally, applying Young’s inequality to the source term yields
+$\operatorname{Re}(f,u_{m})\leq\frac{1}{2}\|f\|^{2}_{L^{2}(0,1)}+\frac{1}{2}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
+Putting all estimates together, we obtain
+$\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}$ $+2\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+2\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}(0,1)}$
+$\leq\|f\|^{2}_{L^{2}(0,1)}+\big(1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\|u_{m}\|^{2}_{L^{2}(0,1)}.$
+We introduce the constant
+$C_{1}:=1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}.$
+Setting $\eta(t):=\|u_{m}(t)\|^{2}_{L^{2}(0,1)}$, the above inequality implies
+$\eta^{\prime}(t)\leq C_{1}\eta(t)+\|f(t)\|^{2}_{L^{2}(0,1)}.$
+By Grönwall’s inequality, for all $t\in[0,T]$,
+$\|u_{m}(t)\|^{2}_{L^{2}(0,1)}\leq e^{C_{1}t}\left(\|u_{0}\|^{2}_{L^{2}(0,1)}+\int_{0}^{t}\|f(s)\|^{2}_{L^{2}(0,1)}\,ds\right).$
+Taking the supremum over $t\in[0,T]$, we conclude
+$\|u_{m}\|^{2}_{L^{\infty}(0,T;L^{2}(0,1))}\leq$ $\;C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
 […]
-$\qquad\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.9)
-where the constant $C>0$ depends only on the final time $T$ and the spatial domain $(0,1)$, and is independent of $m$, the coefficients $a,b,q$, and the data $u_{0}$ and $f$.
-Proof.
-We test the weak form (2.6) with $u_{m}$. This yields
-$Re(\partial_{t}u_{m},u_{m})+Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})+Re(b\,\partial_{x}u_{m},u_{m})+Re(qu_{m},u_{m})=Re(f,u_{m}).$
-The first term is
-$Re(\partial_{t}u_{m},u_{m})=\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
-Using $a\geq\alpha>0$, the second term satisfies
-$Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})\geq\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}.$
-For the third term, integration by parts gives
-$Re(b\,\partial_{x}u_{m},u_{m})=-\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}),$
-so
-$|(b\,\partial_{x}u_{m},u_{m})|\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
-The potential term is
-$Re(qu_{m},u_{m})=Re(\sqrt{q}u_{m},\sqrt{q}u_{m})=\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}.$
-And the right-hand side is bounded by
+where the constant $C$ depends only on $T$.
+Integrating the energy inequality over $[0,T]$, and using the previous bound, we obtain:
+$\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+2\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad\leq C\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\right\}}$
 […]
-Putting all together, we obtain
-$\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}+2\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+2\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}\leq\|f\|^{2}_{L^{2}(0,1)}+C_{1}\|u_{m}\|^{2}_{L^{2}(0,1)},$
-where $C_{1}:=1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}$.
-Letting $\eta(t):=\|u_{m}(t)\|_{L^{2}(0,1)}^{2}$, this gives:
-$\eta^{\prime}(t)\leq C_{1}\eta(t)+\|f(t)\|_{L^{2}(0,1)}^{2}.$
-Applying Grönwall’s inequality, we obtain:
-$\|u_{m}(t)\|_{L^{2}(0,1)}^{2}\leq e^{C_{1}t}\left(\|u_{0}\|_{L^{2}(0,1)}^{2}+\int_{0}^{t}\|f(s)\|_{L^{2}(0,1)}^{2}\,ds\right).$
-Taking the supremum over $t\in[0,T]$, we deduce:
-[…]
-where $C$ depends only on $T$.
-Integrating the energy inequality over $[0,T]$, and using the previous bound, we obtain:
-$\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$
-$\qquad\leq C\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.10)
 where the constant $C$ depends only on $T$.
 To estimate $\partial_{t}u_{m}$, let $\phi\in H^{1}_{0}(0,1)$ with $\|\phi\|_{H^{1}_{0}(0,1)}\leq 1$. Then, from the weak formulation, we have
-$Re\langle\partial_{t}u_{m}(t),\phi\rangle$ $=Re(f(t),\phi)-Re(a(t)\,\partial_{x}u_{m}(t),\partial_{x}\phi)$
-$-Re(b(t)\,\partial_{x}u_{m}(t),\phi)-Re(q\,u_{m}(t),\phi).$
+$\operatorname{Re}\langle\partial_{t}u_{m}(t),\phi\rangle$ $=\operatorname{Re}(f(t),\phi)-\operatorname{Re}(a(t)\,\partial_{x}u_{m}(t),\partial_{x}\phi)$
+$-\operatorname{Re}(b(t)\,\partial_{x}u_{m}(t),\phi)-\operatorname{Re}(q\,u_{m}(t),\phi).$
 Applying Hölder’s inequality and the Poincaré inequality on $(0,1)$, we estimate
@@
 $\|\partial_{t}u_{m}(t)\|_{H^{-1}(0,1)}$ $\leq C_{2}\Big(\|f(t)\|_{L^{2}(0,1)}+\left(\|a(t)\|_{L^{\infty}(0,1)}+\|b(t)\|_{L^{\infty}(0,1)}\right)\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}$
-$+\|q\|^{\frac{1}{2}}_{L^{\infty}}\|\sqrt{q}u_{m}(t)\|_{L^{2}(0,1)}\Big),$
+$+\|q\|_{L^{\infty}(0,1)}\|u_{m}(t)\|_{L^{2}(0,1)}\Big),$
 where
 $C_{2}:=\max\left\{C_{P},1\right\}.$
-Applying (2.2) for $u_{m}$ and $\partial_{x}u_{m}$ by integrating with respect to $t$ over $[0,T]$, we conclude
-$\|\partial_{t}u_{m}\|_{L^{2}(0,T;H^{-1}(0,1))}$ $\leq C\cdot\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}$
-[…]
+We now bound the potential contribution using the previously obtained estimate $u_{m}\in L^{\infty}([0,T];L^{2}(0,1))$. Indeed,
+$\|q\|_{L^{\infty}(0,1)}\|u_{m}(t)\|_{L^{2}(0,1)}\leq\|q\|_{L^{\infty}(0,1)}\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))},$
+and therefore
+$\big\|\,\|q\|_{L^{\infty}(0,1)}\|u_{m}(\cdot)\|_{L^{2}(0,1)}\big\|_{L^{2}(0,T)}\leq\sqrt{T}\,\|q\|_{L^{\infty}(0,1)}\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}.$
+Using (2.2), (2.2) by taking square root and absorbing all numbers into the constant $C$, we conclude
+$\|\partial_{t}u_{m}\|_{L^{2}(0,T;H^{-1}(0,1))}$ $\leq C\cdot\exp{\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}}$
+$\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}([0,T]\times(0,1))}+\|b\|_{L^{\infty}([0,T]\times(0,1))}\right)+\|q\|_{L^{\infty}(0,1)}\right)$
 $\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\right),$ (2.12)
@@
 Theorem 2.3.
-Let $T>0$, and consider the interval $(0,1)\subset\mathbb{R}$. Then, the problem (2.1) admits a unique weak solution
-$u\in L^{2}([0,T];H^{1}_{0}(0,1))\,\,\,\text{with}\,\,\,\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1)),\,\,\,\sqrt{q}u\in L^{2}([0,T];L^{2}(0,1)),$
+Let $T>0$ and consider the interval $(0,1)\subset\mathbb{R}$. Then the problem (2.1) admits a unique weak solution
+$u\in L^{2}([0,T];H_{0}^{1}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),\qquad\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1)),$
+$\sqrt{q_{+}}\,u\in L^{2}([0,T];L^{2}(0,1)),$
 in the sense of Definition 1.
 Proof.
-Let $\{u_{m}\}_{m=1}^{\infty}$ be the Galerkin approximations constructed in (2.4). The uniform energy estimate (2.2) ensures that the following sequences are bounded independently of $m$:
-$\{u_{m}\}\subset L^{2}([0,T];H^{1}_{0}(0,1)),\quad\{\partial_{t}u_{m}\}\subset L^{2}([0,T];H^{-1}(0,1)),$
-$\{\sqrt{q}u_{m}\}\subset L^{2}([0,T];L^{2}(0,1)),\quad\{u_{m}\}\subset L^{\infty}([0,T];L^{2}(0,1)).$
-By Banach–Alaoglu and the Aubin–Lions compactness lemma (see [15, Theorem 8, p. 289]), there exists a subsequence $u_{m_{l}}$ and a limit function $u$ such that:
-$\left\{\begin{aligned} u_{m_{l}}&\rightharpoonup u&&\text{in }L^{2}([0,T];H^{1}_{0}(0,1)),\\ \sqrt{q}u_{m_{l}}&\rightharpoonup\sqrt{q}u&&\text{in }L^{2}([0,T];L^{2}(0,1)),\\ \partial_{t}u_{m_{l}}&\rightharpoonup\partial_{t}u&&\text{in }L^{2}([0,T];H^{-1}(0,1)),\\ u_{m_{l}}&\to u&&\text{in }L^{2}([0,T];L^{2}(0,1)).\end{aligned}\right$.
-(2.12)
-Using these convergence properties and the weak continuity of all terms in the variational formulation, we can pass to the limit in the Galerkin formulation.
-Hence, $u$ satisfies the weak form of (2.1).
-Uniqueness.
-Suppose $u\in L^{2}([0,T];H^{1}_{0}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$, and that $u$ is a weak solution corresponding to zero initial data and zero forcing term, i.e., $u_{0}=0$ and $f=0$.
-Then, for a.e.
-$t\in[0,T]$, testing the weak formulation with $u(t)$ yields:
-$Re\langle\partial_{t}u(t),u(t)\rangle+Re(a\,\partial_{x}u(t),\partial_{x}u(t))+Re(b\,\partial_{x}u(t),u(t))+Re(qu(t),u(t))=0.$
-Using the identity $Re\langle\partial_{t}u(t),u(t)\rangle=\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}$, and integrating by parts in the convective term:
-$Re(b\,\partial_{x}u,u)=-\frac{1}{2}(\partial_{x}b\,,|u|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u\|^{2}_{L^{2}(0,1)}.$
-Together with the assumption $a\geq\alpha>0$ and the positivity of $q$, we obtain the differential inequality:
-$\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}+\alpha\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}+\|\sqrt{q}u(t)\|^{2}_{L^{2}(0,1)}\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u(t)\|^{2}_{L^{2}(0,1)}.$
-Applying Grönwall’s inequality and using $u(0)=0$, we conclude that $u(t)\equiv 0$ for all $t\in[0,T]$, proving uniqueness. ∎
-Proposition 2.4.
-Let $u\in L^{2}([0,T];H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ be the weak solution to the parabolic problem (2.1), and assume the coefficients satisfy:
-$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
-Then, the weak solution to the parabolic equation (2.1) we have the following estimates:
-$\Big\{\alpha^{1/2}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))};\left.\;\|u\|_{L^{\infty}([0,T];L^{2}(0,1))};\;\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}\right\}$
-$\qquad\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.13)
-$\|\partial_{t}u\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
+Let $\{u_{m}\}_{m=1}^{\infty}$ be the Galerkin approximations constructed in (2.4). The uniform estimate (2.2) implies that, uniformly in $m$,
+  • $u_{m}\ \text{is bounded in }L^{2}(0,T;H_{0}^{1}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),$
+  • $\partial_{t}u_{m}\ \text{is bounded in }L^{2}([0,T];H^{-1}(0,1)),$
+  • $\sqrt{q_{+}}\,u_{m}\ \text{is bounded in }L^{2}([0,T];L^{2}(0,1)).$
+By Banach–Alaoglu and the Aubin–Lions lemma (see [18, Theorem 8, p. 289]), there exist a subsequence $(u_{m_{l}})$ and a function $u$ such that
 […]
-$\qquad\times\left(\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|u_{0}\|_{L^{2}(0,1)}\right),$ (2.14)
-$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2})};\,\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2})};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})};\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2})}\Big\}$
-$\quad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
-$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\quad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.15)
-where the constant $C>0$ depends only on the final time $T$ and the spatial interval $(0,1)$, but not on the data or coefficients.
+Using these convergences and the boundedness of $a,b$ on $0,T]\times(0,1)$ and of $q$ on $(0,1)$, we can pass to the limit in (2.6) and conclude that $u$ satisfies (1).
+Moreover, by the standard Lions lemma (see [23, Lemma 1.2, p.
+260]), $u\in L^{\infty}([0,T];L^{2}(0,1))$ and $u(0)=u_{0}$.
+Uniqueness. Let $u$ be a weak solution with $u_{0}=0$ and $f=0$. Testing the weak formulation with $u(t)$ and taking real parts yields
+$\operatorname{Re}\langle\partial_{t}u(t),u(t)\rangle+\operatorname{Re}(a\,\partial_{x}u(t),\partial_{x}u(t))+\operatorname{Re}(b\,\partial_{x}u(t),u(t))+\operatorname{Re}(qu(t),u(t))=0.$
+Since $u\in L^{2}([0,T];H_{0}^{1})$ and $\partial_{t}u\in L^{2}([0,T];H^{-1})$, we have the identity
+$\operatorname{Re}\langle\partial_{t}u(t),u(t)\rangle=\frac{1}{2}\frac{d}{dt}\|u(t)\|_{L^{2}(0,1)}^{2}\quad\text{in }\mathcal{D}^{\prime}(0,T).$
+Moreover, integrating by parts gives
+$\operatorname{Re}(b\,\partial_{x}u,u)=-\frac{1}{2}(\partial_{x}b,|u|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|u\|_{L^{2}(0,1)}^{2}.$
+Writing $q=q_{+}-q_{-}$ with $q_{+}\geq 0$ and $q_{-}:=\max\{-q,0\}$, we obtain
+$\operatorname{Re}(qu,u)=(q_{+},|u|^{2})-(q_{-},|u|^{2})\geq\|\sqrt{q_{+}}\,u\|_{L^{2}(0,1)}^{2}-\|q_{-}\|_{L^{\infty}(0,1)}\|u\|_{L^{2}(0,1)}^{2}.$
+Using also $\operatorname{Re}(a\partial_{x}u,\partial_{x}u)\geq\alpha\|\partial_{x}u\|_{L^{2}(0,1)}^{2}$, we arrive at
+$\frac{1}{2}\frac{d}{dt}\|u(t)\|_{L^{2}(0,1)}^{2}+\alpha\|\partial_{x}u(t)\|_{L^{2}(0,1)}^{2}+\|\sqrt{q_{+}}\,u(t)\|_{L^{2}(0,1)}^{2}\leq C\,\|u(t)\|_{L^{2}(0,1)}^{2},$
+where $C:=\tfrac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{-}\|_{L^{\infty}(0,1)}$. Since $u(0)=0$, Grönwall’s inequality implies $u\equiv 0$ on $[0,T]$. ∎
+Lemma 2.4 (Higher time–space regularity).
+Let $u$ be the unique weak solution of (2.1) given by Theorem 2.3 with initial data $u_{0}\in H^{1}_{0}(0,1)$ and $f\in L^{2}(0,T;L^{2}(0,1))$. Assume that the coefficients satisfy
+$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}(0,T;L^{\infty}(0,1)),\qquad a(t,x)\geq\alpha>0,$
+and
+$q\in L^{\infty}(0,1),\qquad q=q_{+}-q_{-},\quad q_{+}\geq 0,\quad q_{-}:=\max\{-q,0\}.$
+Then the solution $u$ satisfies:
+$\|u\|_{L^{\infty}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q_{+}}\,u\|_{L^{2}(0,T;L^{2}(0,1))}$
+$\qquad\leq C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
+$\qquad\quad\Big(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.14)
+$\|\partial_{t}u\|_{L^{2}(0,T;H^{-1}(0,1))}\leq C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
+$\qquad\times\left(1+\frac{1}{\sqrt{\alpha}}\Big(\|a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\Big)+\|q\|_{L^{\infty}(0,1)}\right)$
+$\qquad\times\Big(\|f\|_{L^{2}(0,T;L^{2}(0,1))}+\|u_{0}\|_{L^{2}(0,1)}\Big),$ (2.15)
+and moreover
+$\|\partial_{t}u\|_{L^{2}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u\|_{L^{\infty}(0,T;L^{2}(0,1))}+\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2}(0,1))}$
+$\quad+\|\partial_{x}^{2}u\|_{L^{2}(0,T;L^{2}(0,1))}\leq\frac{C}{\alpha}\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
+$\qquad\times\Big(1+\|q\|_{L^{\infty}(0,1)}+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}$
+$\qquad+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)\times\Big(\sqrt{\alpha}\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.16)
+where $C>0$ depends only on $T$ and the interval $(0,1)$.
 Proof.
-By taking limits in (2.2)-(2.2) as $m=m_{l}\to\infty$, we derive the estimates (2.4)-(2.4).
-To prove (2.4), we start from the equation in strong form:
-$\partial_{t}u-\partial_{x}(a\partial_{x}u)=-b\partial_{x}u-qu+f.$
-Multiply both sides by $\partial_{t}u$ and integrate over $(0,1)$:
-$\|\partial_{t}u\|^{2}_{L^{2}}$ $-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)+Re(b\partial_{x}u\,,\partial_{t}u)$
-$+Re(qu\,,\partial_{t}u)=Re(f\,,\partial_{t}u).$
-We handle the second term by integration by parts:
-$-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)=Re(a\partial_{x}u\,,\partial_{x}(\partial_{t}u))=\frac{1}{2}\frac{d}{dt}(a\,,|\partial_{x}u|^{2})-\frac{1}{2}(\partial_{t}a\,,|\partial_{x}u|^{2})$
-$\qquad\geq\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}-\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)},$
-where the second term we move to the right hand-side.
-Apply Cauchy–Schwarz and Young’s inequality to the remaining terms:
-$\left|(b\partial_{x}u\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}},$
-$\left|(qu\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}},$
-$\left|(f\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
-Combining everything, we obtain:
-$\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}$ $\leq\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}$
-$+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
-Now integrate over $[0,T]$:
-$\frac{1}{4}\|\partial_{t}u(t)\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u\|^{2}_{L^{\infty}([0,T];L^{2}(0,1))}$
-$\qquad\leq\left(\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|^{2}_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\|\partial_{x}u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|q\|^{2}_{L^{\infty}(0,1)}\|u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\|f\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u(0)\|^{2}_{L^{2}(0,1)}.$
-We now estimate the last term $\|\partial_{x}u(0)\|_{L^{2}}$ by using initial condition, we obtain
-$\|\partial_{x}u(0)\|_{L^{2}(0,1)}=\|\partial_{x}u_{0}\|_{L^{2}(0,1)}=\|u_{0}\|_{H^{1}_{0}(0,1)}.$
-Using the full energy estimate, we also have bounds for $\|u\|_{L^{2}([0,T];L^{2})}$, $\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}$ in terms of data. Therefore:
-$\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))}+\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
-$\qquad\times\left(\|a\|^{1/2}_{L^{\infty}}+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|_{L^{\infty}}\right)$
-$\qquad\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|^{\frac{1}{2}}_{L^{\infty}}\right)\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$
-where $C$ depends only on $T$ and the domain $(0,1)$.
-Now, use the PDE again to write:
-$\partial_{x}(a\partial_{x}u)=\partial_{t}u+b\partial_{x}u+qu-f.$
-Then
-$\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})}$ $\leq\|\partial_{t}u\|_{L^{2}}+\|b\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}}+\|q\|_{L^{\infty}}\|u\|_{L^{2}}+\|f\|_{L^{2}},$
-which again is controlled by data via the energy estimate.
-Finally, to estimate the second spatial derivative $\partial_{x}^{2}u$, we use the identity
-$a\,\partial_{x}^{2}u=\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u,$
-and the ellipticity condition $a(t,x)\geq\alpha>0$. Then, by the triangle inequality and Hölder’s inequality:
-$\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$ $=\left\|\frac{1}{a}\left(\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u\right)\right\|_{L^{2}}$
-$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\,\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right)$
-$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\|_{L^{\infty}([0,T]\times(0,1))}\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right).$
-Therefore, $\partial_{x}^{2}u\in L^{2}([0,T];L^{2}(0,1))$ with a constant depending only on $T$ and the interval $(0,1)$.
-This completes the proof, with all terms estimated in terms of data and coefficients, and the constant depending only on $T$ and the interval $(0,1)$. ∎
-3. Non-homogeneous Dirichlet condition
-We now consider the parabolic equation in divergence form (1) with diffusion coefficient, drift term, and potential with initial condition (1.2) and non-homogeneous Dirichlet boundary conditions (1.3).
-We decompose the solution as
-$u(t,x)=w(t,x)+\psi(t,x),$ (3.1)
-where $w$ satisfies homogeneous Dirichlet boundary conditions
-$w(t,0)=w(t,1)=0,$
-and $\psi$ is a lifting function defined by
-$\psi(t,x):=(1-x)g_{0}(t)+xg_{1}(t),\quad(t,x)\in[0,T]\times(0,1).$
-This function satisfies:
-  • $\psi(t,0)=g_{0}(t),\quad\psi(t,1)=g_{1}(t)$,
-  • $\partial_{x}\psi(t,x)=g_{1}(t)-g_{0}(t)$ is independent of $x$,
-  • $\psi\in H^{1}([0,T];H^{1}(0,1))$ provided $g_{0},g_{1}\in H^{1}([0,T])$.
-Substituting (3.1) into (1), we obtain the equation for $w$:
-$\left\{\begin{array}[]{l}\partial_{t}w-\partial_{x}(a(t,x)\partial_{x}w)+b(t,x)\partial_{x}w+q(x)w=\tilde{f}(t,x),\\[2.0pt] w(0,x)=u_{0}(x)-\psi(0,x),\quad x\in(0,1),\\[2.0pt] w(t,0)=w(t,1)=0,\quad t\in[0,T],\end{array}\right$.
-(3.2)
-where the inhomogeneous source term is given by
-$\tilde{f}(t,x)=$ $f(t,x)-\partial_{t}\psi(t,x)+\partial_{x}(a(t,x)\partial_{x}\psi(t,x))-b(t,x)\partial_{x}\psi(t,x)-q(x)\psi(t,x)$
-$=$ $f(t,x)-(1-x)g_{0}^{\prime}(t)-xg_{1}^{\prime}(t)+(\partial_{x}a(t,x)-b(t,x))(g_{1}(t)-g_{0}(t))$
-$-q(x)\big((1-x)g_{0}(t)+xg_{1}(t)\big).$
-The right-hand side $\tilde{f}\in L^{2}([0,T];L^{2}(0,1))$ under the assumptions $f\in L^{2}$ and $g_{0},g_{1}\in H^{1}([0,T])$, and the fact that the coefficients are bounded.
-Moreover, $\|\tilde{f}\|_{L^{2}}$ can be bounded by
-$\|\tilde{f}\|_{L^{2}}\leq\|f\|_{L^{2}}+C\left(\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
-for some constant $C>0$ depending only on $\|a\|_{W^{1,\infty}},\|b\|_{L^{\infty}},\|q\|_{L^{\infty}}$.
-We now state the energy estimate.
-Theorem 3.1.
-Let $u$ solve the parabolic problem (1)–(1.3), with initial data $u_{0}\in L^{2}(0,1)$, source term $f\in L^{2}([0,T];L^{2}(0,1))$, and boundary data $g_{0},g_{1}\in H^{1}([0,T])$.
-Assume the coefficients satisfy
-$a(t,x)\geq\alpha>0,\quad a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$
-Then the solution $u\in L^{2}([0,T];H^{1}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$ satisfies:
-$\|u\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}$
-[…]
-$\qquad\times\Big(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\Big),$ (3.3)
-where $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
-Proof.
-As described above, we write $u=w+\psi$, where $w$ satisfies the homogeneous problem (3.2). Applying the energy estimate for $w$ from Theorem 2.2, we obtain:
-$\|w\|_{L^{\infty}L^{2}}+\sqrt{\alpha}\|\partial_{x}w\|_{L^{2}L^{2}}+\|\sqrt{q}w\|_{L^{2}L^{2}}\leq Ce^{T\|\partial_{x}b\|_{L^{\infty}}}\left(\|u_{0}-\psi(0)\|_{L^{2}}+\|\tilde{f}\|_{L^{2}}\right).$ (3.4)
-Since $\psi(0,x)=(1-x)g_{0}(0)+xg_{1}(0)$, we estimate:
-$\|\psi(0)\|_{L^{2}}\leq C\left(|g_{0}(0)|+|g_{1}(0)|\right)\leq C\left(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}\right),$
-hence,
-$\|u_{0}-\psi(0)\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}+C(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
-Next, since $u=w+\psi$, we use the triangle inequality:
-$\|u\|_{L^{\infty}L^{2}}$ $\leq\|w\|_{L^{\infty}L^{2}}+\|\psi\|_{L^{\infty}L^{2}},$
-$\|\partial_{x}u\|_{L^{2}L^{2}}$ $\leq\|\partial_{x}w\|_{L^{2}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}},$
-$\|\sqrt{q}u\|_{L^{2}L^{2}}$ $\leq\|\sqrt{q}w\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}.$
-Note that $\psi(t,x)\in\text{span}\{1,x\}\cdot g_{j}(t)$, so
-$\|\psi\|_{L^{\infty}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}\leq C(T)\left(1+\|a\|^{1/2}_{L^{\infty}}+\|q\|^{1/2}_{L^{\infty}}\right)(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
-Combining this with (3.4) yields (3.1), which completes the proof of Theorem 3.1. ∎
-Corollary 3.2.
-Let $u$ solve the non-homogeneous Dirichlet problem (1)–(1.3) with data
-$f\in L^{2}([0,T];L^{2}(0,1)),\quad u_{0}\in L^{2}(0,1),\quad g_{0},g_{1}\in H^{1}([0,T]),$
-and assume the coefficients satisfy:
-$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
-Then:
-$\partial_{t}u,\,\partial_{x}(a\partial_{x}u)\in L^{2}([0,T];L^{2}(0,1)),$
-and the following estimate holds:
-$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial^{2}_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}\Big\}$
-$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
-$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\qquad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
-where the constant $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
-Proof.
-We decompose the solution as $u=w+\psi$, where $w$ solves the corresponding homogeneous Dirichlet problem with modified source
-$\tilde{f}=f-\mathcal{L}\psi,\qquad\psi(t,x)=(1-x)g_{0}(t)+xg_{1}(t)$
-is the boundary lifting.
-Then, since $\partial_{t}\psi\in L^{2}([0,T];L^{2}(0,1))$ and $\partial_{x}(a\partial_{x}\psi)\in L^{2}([0,T];L^{2}(0,1))$, and the estimate from Proposition 2.4 applies to $w$, we have:
-$\|\partial_{t}u\|_{L^{2}}\leq\|\partial_{t}w\|_{L^{2}}+\|\partial_{t}\psi\|_{L^{2}},\quad\|\partial_{x}(a\partial_{x}u)\|_{L^{2}}\leq\|\partial_{x}(a\partial_{x}w)\|_{L^{2}}+\|\partial_{x}(a\partial_{x}\psi)\|_{L^{2}}.$
-Now, note that $\psi\in H^{1}([0,T];H^{1}(0,1))$, and all terms involving $\psi$ can be bounded in terms of $\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}$.
-Combining with the bound for $w$ from Proposition 2.4 yields the desired result.
-∎
-4. Very weak solutions
+The estimates (2.4)–(2.4) follow by passing to the limit in the Galerkin estimates (2.2).
+We prove (2.4) at the Galerkin level and then pass to the limit.
+Let $u_{m}$ be the Galerkin approximations.
+Since $u_{m}\in C^{1}([0,T];V_{m})$, all computations below are justified.
+From the Galerkin equation we have (in the finite-dimensional sense)
+$\partial_{t}u_{m}-\partial_{x}(a\partial_{x}u_{m})=-b\,\partial_{x}u_{m}-q\,u_{m}+f.$
+Multiply by $\partial_{t}u_{m}$ and integrate over $(0,1)$:
+$\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}$- $\operatorname{Re}\big(\partial_{x}(a\partial_{x}u_{m}),\partial_{t}u_{m}\big)$
+$=-\operatorname{Re}\big(b\,\partial_{x}u_{m},\partial_{t}u_{m}\big)-\operatorname{Re}\big(q\,u_{m},\partial_{t}u_{m}\big)+\operatorname{Re}\big(f,\partial_{t}u_{m}\big).$
+Integrating by parts,
+$-\operatorname{Re}\big(\partial_{x}(a\partial_{x}u_{m}),\partial_{t}u_{m}\big)$ $=\operatorname{Re}\big(a\partial_{x}u_{m},\partial_{x}\partial_{t}u_{m}\big)$
+$=\frac{1}{2}\frac{d}{dt}\,(a,|\partial_{x}u_{m}|^{2})-\frac{1}{2}(\partial_{t}a,|\partial_{x}u_{m}|^{2})$
+$\geq\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}-\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\,\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}.$
+Using Cauchy–Schwarz and Young’s inequality,
+$\big|\big(b\partial_{x}u_{m},\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}^{2}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2},$
+$\big|\big(qu_{m},\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|q\|_{L^{\infty}(0,1)}^{2}\|u_{m}\|_{L^{2}(0,1)}^{2},$
+$\big|\big(f,\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|f\|_{L^{2}(0,1)}^{2}.$
+Combining these estimates yields
+$\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}$ $\leq\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2}+\|b\|_{L^{\infty}}^{2}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2}$
+$\quad+\|q\|_{L^{\infty}}^{2}\|u_{m}\|_{L^{2}(0,1)}^{2}+\|f\|_{L^{2}(0,1)}^{2}.$
+Integrating the above inequality over $(0,T)$ and taking the supremum with respect to $t\in[0,T]$, we obtain
+$\frac{1}{4}\|\partial_{t}u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u_{m}\|^{2}_{L^{\infty}(0,T;L^{2}(0,1))}$
+$\qquad\leq\left(\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|^{2}_{L^{\infty}(0,T;L^{\infty}(0,1))}\right)\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}$
+$\qquad\quad+\|q\|^{2}_{L^{\infty}(0,1)}\|u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\|f\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u_{m}(0)\|^{2}_{L^{2}(0,1)}.$
+Using the initial condition $u_{m}(0)=P_{m}u_{0}$ and the stability of the projection $P_{m}$ in $H^{1}_{0}(0,1)$, we have
+$\|\partial_{x}u_{m}(0)\|_{L^{2}(0,1)}\leq\|u_{0}\|_{H^{1}_{0}(0,1)}.$
+Moreover, the energy estimate (2.4) yields bounds for $\|u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$ and
+$\|\partial_{x}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$ in terms of the data. Consequently, there exists a constant $C>0$, independent of $m$, such that
+$\|\partial_{t}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u_{m}\|_{L^{\infty}(0,T;L^{2}(0,1))}$
+$\quad\leq C\exp\!\Big\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\Big\}$
+$\qquad\times\Big(1+\|q\|_{L^{\infty}(0,1)}+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
+$\qquad\times\Big(\sqrt{\alpha}\,\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.17)
+where $C$ depends only on $T$.
+The previous estimate is uniform in $m$, hence $\{\partial_{t}u_{m}\}$ is bounded in $L^{2}(0,T;L^{2}(0,1))$.
+By Banach–Alaoglu, there exist a subsequence (not relabelled) and a function $w\in L^{2}(0,T;L^{2}(0,1))$ such that
+$\partial_{t}u_{m}\rightharpoonup w\qquad\text{in }L^{2}(0,T;L^{2}(0,1)).$
+On the other hand, from the energy estimate (2.4) we already know that $\{\partial_{t}u_{m}\}$ is bounded in $L^{2}(0,T;H^{-1}(0,1))$ and, along the same subsequence,
+$\partial_{t}u_{m}\rightharpoonup\partial_{t}u\qquad\text{in }L^{2}(0,T;H^{-1}(0,1)).$
+Since the embedding $L^{2}(0,1)\hookrightarrow H^{-1}(0,1)$ is continuous, the weak limit in
+$L^{2}(0,T;L^{2}(0,1))$ is also the weak limit in $L^{2}(0,T;H^{-1}(0,1))$. Therefore $w=\partial_{t}u$. In particular,
+$\partial_{t}u\in L^{2}(0,T;L^{2}(0,1))\quad\text{and}\quad\|\partial_{t}u\|_{L^{2}(0,T;L^{2}(0,1))}\leq\liminf_{m\to\infty}\|\partial_{t}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}.$
+Similarly, by weak-* compactness in $L^{\infty}(0,T;L^{2}(0,1))$ we have (up to a subsequence)
+$\partial_{x}u_{m}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\partial_{x}u\qquad\text{in }L^{\infty}(0,T;L^{2}(0,1)),$
+and hence
+$\|\partial_{x}u\|_{L^{\infty}(0,T;L^{2}(0,1))}\leq\liminf_{m\to\infty}\|\partial_{x}u_{m}\|_{L^{\infty}(0,T;L^{2}(0,1))},$
+by weak-* lower semicontinuity of the $L^{\infty}$-norm. Consequently, the first two terms in (2.4) pass to the limit and the same bound holds for $u$.
+Next, from the equation we write
+$\partial_{x}(a\partial_{x}u)=\partial_{t}u+b\,\partial_{x}u+q\,u-f,$
+so
+$\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2})}\leq$ $\|\partial_{t}u\|_{L^{2}(0,T;L^{2})}+\|b\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}(0,T;L^{2})}$
+$+\|q\|_{L^{\infty}}\|u\|_{L^{2}(0,T;L^{2})}+\|f\|_{L^{2}(0,T;L^{2})},$
+which is controlled by the previous bounds and (2.4). Finally,
+$a\,\partial_{x}^{2}u=\partial_{x}(a\partial_{x}u)-(\partial_{x}a)\,\partial_{x}u,$
+hence using $a\geq\alpha$,
+$\|\partial_{x}^{2}u\|_{L^{2}(0,T;L^{2})}\leq\frac{1}{\alpha}\Big(\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2})}+\|\partial_{x}a\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}(0,T;L^{2})}\Big),$
+and (2.4) follows by collecting the estimates and absorbing numerical factors into $C$. ∎
+3. Very weak solutions for the non-homogeneous Dirichlet conditions
 In this section, we consider the parabolic equation on $0,T]\times(0,1)$, where the diffusion coefficient, drift term, potential, initial data, and boundary data are modelled by distributions (e.g., Dirac-type functions).
@@
     $\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\leq C\varepsilon^{-N}.$
+We additionally assume that the negative part $q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}$ satisfies
+$\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\leq C|\log\varepsilon|.$
 Definition 3 (Coefficient moderateness).
@@
     The net $(b_{\varepsilon})\subset L^{\infty}([0,T];W^{1,\infty}(0,1))$ is log-$L^{\infty}(W^{1,\infty})$-moderate if
-    $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\leq C|\log\varepsilon|^{N},\quad\|b_{\varepsilon}\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\leq C\varepsilon^{-N}.$
+    $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\leq C|\log\varepsilon|,\quad\|b_{\varepsilon}\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\leq C\varepsilon^{-N}.$
 Definition 4 (Data moderation).
@@
 Remark 3.1.
-The above assumptions are natural in the context of distributional coefficients, especially in one spatial dimension.
-Regularisations of distributions often form moderate nets.
-In particular, by the structure theorem for distributions (see, e.g., [16]), we have the inclusion
-$\mathcal{D}^{\prime}(0,1)\subset\{\text{$L^{\infty}(0,1)$-moderate families}\}.$ (4.1)
-This means that even if a solution to a boundary value problem on $(0,1)$ does not exist in the classical distributional sense, it may still exist as a net of $L^{\infty}$-moderate functions.
-This framework allows for a consistent treatment of singular coefficients and data.
+The above moderateness assumptions are natural in the context of distributional coefficients and data, particularly in one spatial dimension.
+Let $v\in\mathcal{D}^{\prime}(0,1)$.
+After extension by zero to $\mathbb{R}$, the standard mollification $v_{\varepsilon}:=v*\psi_{\varepsilon}$ defines a net of smooth functions.
+By the structure theorem for distributions (see, e.g., [19]), there exist constants $C>0$ and $N\in\mathbb{N}$, depending on the order of $v$, such that
+$\|v_{\varepsilon}\|_{L^{\infty}(0,1)}\leq C\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
+Hence every distribution admits an $L^{\infty}$-moderate regularisation.
+This observation justifies the use of moderate nets to represent distributional coefficients and data, even in situations where a classical distributional solution of the boundary value problem does not exist.
 As an example, let $f\in L^{2}(0,1)$, where $f:(0,1)\to\mathbb{C}$. We define the extension of $f$ by zero to all of $\mathbb{R}$ as:
-$\widetilde{f}(t,x):=\begin{cases}f(x),&(x)\in(0,1),\\ 0,&(x)\in\mathbb{R}\setminus(0,1).\end{cases}$
+$\widetilde{f}(x):=\begin{cases}f(x),&(x)\in(0,1),\\ 0,&(x)\in\mathbb{R}\setminus(0,1).\end{cases}$
 […]
@@
 Then:
-  • Each $\widetilde{f}_{\varepsilon}\in C^{\infty}(\mathbb{R})$,
-  • The net $(\widetilde{f}_{\varepsilon})$ is $L^{p}(\mathbb{R})$-moderate for any $1\leq p\leq\infty$,
-  • On the original domain, $\widetilde{f}_{\varepsilon}\to f$ in $L^{p}(0,1)$ as $\varepsilon\to 0$,
+  • each $\widetilde{f}_{\varepsilon}\in C^{\infty}(\mathbb{R})$;
+  • for each fixed $2\leq p\leq\infty$, the net $(\widetilde{f}_{\varepsilon})$ is $L^{p}(\mathbb{R})$-moderate;
+  • on the original domain, $\widetilde{f}_{\varepsilon}\to f$ in $L^{p}(0,1)$ as $\varepsilon\to 0$;
   • outside $(0,1)$, the mollified functions decay, i.e., $\|\widetilde{f}_{\varepsilon}\|_{L^{p}(\mathbb{R}\setminus(0,1))}\to 0$.
@@
 (3.1)
-Definition 6.
-Assume that $a\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, and $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$.
-A net of functions $(u_{\varepsilon})_{\varepsilon>0}$ is said to be a very weak solution to the initial/boundary value problem (4.2), if there exist an $L^{\infty}$-moderate regularisation $q_{\varepsilon}$ of $q$, a $W^{1,\infty}$-moderate regularisation $a_{\varepsilon}$ of $a$, a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation $b_{\varepsilon}$ of $b$, an $H^{1}$-moderate regularisation $u_{0,\varepsilon}$ of $u_{0}$, time-distributional moderate regularisations $g_{0,\varepsilon}$ and $g_{1,\varepsilon}$ of $g_{0}$ and $g_{1}$ respectively, and an $L^{2}$-moderate regularisation $f_{\varepsilon}$ of $f$, such that the regularised problem
-$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),\\ \hfill(t,x)\in[0,T]\times(0,1),\\ u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\quad x\in(0,1),\\ u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\quad t\in[0,T],\end{array}\right$.
+Definition 6 (Very weak solution).
+Assume that
+$a\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\qquad b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\qquad q\in\mathcal{D}^{\prime}(0,1),$
+$f\in\mathcal{D}^{\prime}([0,T]\times(0,1)),\qquad u_{0}\in\mathcal{D}^{\prime}(0,1),\qquad g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T]).$
+A net of functions $(u_{\varepsilon})_{\varepsilon\in(0,1]}$ is said to be a very weak solution to the initial–boundary value problem (3.1) if there exist
+  • an $L^{\infty}$-moderate regularisation $q_{\varepsilon}$ of $q$,
+  • a $W^{1,\infty}$-moderate regularisation $a_{\varepsilon}$ of $a$,
+  • a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation $b_{\varepsilon}$ of $b$,
+  • an $H^{1}$-moderate regularisation $u_{0,\varepsilon}$ of $u_{0}$,
+  • time-distributional moderate regularisations $g_{0,\varepsilon}$, $g_{1,\varepsilon}$ of $g_{0}$, $g_{1}$, respectively,
+  • an $L^{2}$-moderate regularisation $f_{\varepsilon}$ of $f$,
+such that, for every $\varepsilon\in(0,1]$, the function $u_{\varepsilon}$ is the unique weak solution of the regularised problem
+$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}\!\big(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)\big)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),\\[2.0pt] \hfill(t,x)\in[0,T]\times(0,1),\\[2.0pt] u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\qquad x\in(0,1),\\[2.0pt] u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\qquad t\in[0,T],\end{array}\right$.
 (3.2)
-holds for all $\varepsilon\in(0,1]$, and the net $(u_{\varepsilon})$ is moderate in the sense that there exist constants $C>0$, $N\in\mathbb{N}$ such that
-$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},\quad\text{for all }\varepsilon\in(0,1].$
+where the regularised diffusion coefficient satisfies the uniform ellipticity condition
+$a_{\varepsilon}(t,x)\geq\alpha>0,\qquad(t,x)\in[0,T]\times(0,1),\ \varepsilon\in(0,1].$
+Moreover, the net $(u_{\varepsilon})$ is required to be moderate in the sense that there exist constants $C>0$ and $N\in\mathbb{N}$ such that
+$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\,\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
+In the previous section, we considered the parabolic problem with regular coefficients and homogeneous Dirichlet boundary conditions.
+Here, we admit distributional coefficients and data, and, after regularisation, handle the non-homogeneous Dirichlet problem by homogenising the boundary conditions via a lifting.
+We decompose the solution as
+$u_{\varepsilon}(t,x)=w_{\varepsilon}(t,x)+\psi_{\varepsilon}(t,x),$ (3.3)
+where $w_{\varepsilon}$ satisfies homogeneous Dirichlet boundary conditions
+$w_{\varepsilon}(t,0)=w_{\varepsilon}(t,1)=0,$
+and $\psi_{\varepsilon}$ is a lifting function defined by
+$\psi_{\varepsilon}(t,x):=(1-x)g_{0,\varepsilon}(t)+xg_{1,\varepsilon}(t),\quad(t,x)\in[0,T]\times(0,1).$
+This function satisfies:
+  • $\psi_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad\psi_{\varepsilon}(t,1)=g_{1,\varepsilon}(t)$,
+  • $\partial_{x}\psi_{\varepsilon}(t,x)=g_{1,\varepsilon}(t)-g_{0,\varepsilon}(t)$ is independent of $x$,
+  • $\psi_{\varepsilon}\in H^{1}([0,T];H^{1}(0,1))$ provided $g_{0,\varepsilon},g_{1,\varepsilon}\in H^{1}([0,T])$.
+Substituting (3.3) into (3.2), we obtain the equation for $w_{\varepsilon}$:
+$\left\{\begin{array}[]{l}\partial_{t}w_{\varepsilon}-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}w_{\varepsilon})+b_{\varepsilon}(t,x)\partial_{x}w_{\varepsilon}+q_{\varepsilon}(x)w_{\varepsilon}=\widetilde{f}_{\varepsilon}(t,x),\\[2.0pt] w_{\varepsilon}(0,x)=u_{0,\varepsilon}(x)-\psi_{\varepsilon}(0,x),\quad x\in(0,1),\\[2.0pt] w_{\varepsilon}(t,0)=w_{\varepsilon}(t,1)=0,\quad t\in[0,T],\end{array}\right$.
+(3.4)
+where the inhomogeneous source term is given by
+$\widetilde{f}_{\varepsilon}(t,x)=$ $f_{\varepsilon}(t,x)-\partial_{t}\psi_{\varepsilon}(t,x)+\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x))-b_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x)$
+$-q_{\varepsilon}(x)\psi_{\varepsilon}(t,x)=f_{\varepsilon}(t,x)-(1-x)g_{0,\varepsilon}^{\prime}(t)-xg_{1,\varepsilon}^{\prime}(t)$
+$+(\partial_{x}a_{\varepsilon}(t,x)-b_{\varepsilon}(t,x))(g_{1,\varepsilon}(t)-g_{0,\varepsilon}(t))-q_{\varepsilon}(x)\big((1-x)g_{0,\varepsilon}(t)+xg_{1,\varepsilon}(t)\big),$
+where $\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x))=(\partial_{x}a_{\varepsilon})(g_{1,\varepsilon}-g_{0,\varepsilon})$ since $\partial^{2}_{x}\psi_{\varepsilon}=0$.
+The right-hand side $\widetilde{f}_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1))$ under the assumptions
+$f_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1)),\quad a_{\varepsilon},\,\partial_{x}a_{\varepsilon},\,b_{\varepsilon}\in L^{\infty}([0,T];L^{\infty}(0,1))\quad q_{\varepsilon}\in L^{\infty}(0,1)$
+and $g_{0,\varepsilon},g_{1,\varepsilon}\in H^{1}([0,T])$, and the fact that the coefficients are bounded. Moreover, using $H^{1}(0,T)\hookrightarrow L^{\infty}(0,T)$, we obtain
+$\|\widetilde{f}_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\leq\;$ $\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g^{\prime}_{0,\varepsilon}\|_{L^{2}(0,T)}+\|g^{\prime}_{1,\varepsilon}\|_{L^{2}(0,T)}$
+$+\Big(\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T]\times(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T]\times(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\Big)$
+$\qquad\times\Big(\|g_{0,\varepsilon}\|_{L^{\infty}(0,T)}+\|g_{1,\varepsilon}\|_{L^{\infty}(0,T)}\Big)$
+$\leq\;$ $\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+C\Big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}}+\|b_{\varepsilon}\|_{L^{\infty}}+\|q_{\varepsilon}\|_{L^{\infty}}\Big)$
+$\qquad\times\Big(\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\Big),$ (3.5)
+where the constant $C$ depends only on $T$.
+Estimate of the initial data.
+Since
+$\psi_{\varepsilon}(0,x)=(1-x)g_{0,\varepsilon}(0)+xg_{1,\varepsilon}(0),$
+we have by the triangle inequality
+$\|w_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}=\|u_{0,\varepsilon}-\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}.$
+Moreover,
+$\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}\big|(1-x)g_{0,\varepsilon}(0)+xg_{1,\varepsilon}(0)\big|^{2}\,dx\leq C\big(|g_{0,\varepsilon}(0)|^{2}+|g_{1,\varepsilon}(0)|^{2}\big),$
+hence
+$\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq C\big(|g_{0,\varepsilon}(0)|+|g_{1,\varepsilon}(0)|\big).$
+Finally, by the trace inequality $|h(0)|\leq C\|h\|_{H^{1}(0,T)}$ (valid for all $h\in H^{1}(0,T)$), we obtain
+$\|w_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+C\big(\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\big),$ (3.6)
+where $C>0$ depends only on $T$ and the interval $(0,1)$.
+Writing $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, the function $w_{\varepsilon}$ is the unique weak solution of the homogeneous problem (3.4) with initial datum $u_{0,\varepsilon}-\psi_{\varepsilon}(0)$ and source term $\widetilde{f}_{\varepsilon}$.
+Proposition 3.1.
+Let $u_{\varepsilon}$ be the weak solution to (3.2) with $u_{0,\varepsilon}\in L^{2}(0,1)$ and $f_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$. Assume
+$a_{\varepsilon}(t,x)\geq\alpha>0,\qquad a_{\varepsilon}\in L^{\infty}(0,T;L^{\infty}(0,1)),\qquad b_{\varepsilon}\in L^{\infty}(0,T;W^{1,\infty}(0,1)),$
+and
+$q_{\varepsilon}\in L^{\infty}(0,1),\qquad q_{\varepsilon}=q_{+,\varepsilon}-q_{-,\varepsilon},\quad q_{+,\varepsilon}\geq 0,\quad q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}.$
+Then $u_{\varepsilon}\in L^{2}(0,T;H^{1}_{0}(0,1))$ and $\partial_{t}u_{\varepsilon}\in L^{2}(0,T;H^{-1}(0,1))$, and
+$\|u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\,u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}$
+$\qquad\leq C\Big(\exp\!\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}$
+$\qquad\quad\times\big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\big)$
+$\qquad\quad\times\Big(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\Big)$
+$\qquad\quad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big),$ (3.7)
+where the constant $C$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
+The estimate (3.1) shows that $u_{\varepsilon}$ is a moderate net whenever the regularisations of the coefficients and data are moderate in the sense of Definition 6.
+Proof.
+We write $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, where $w_{\varepsilon}$ satisfies the homogeneous problem (3.4). Applying the energy estimate from Theorem 2.2, we obtain
+$\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad\leq C\exp{\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}}$
+$\qquad\quad\times\big(\|u_{0,\varepsilon}-\psi_{\varepsilon}(0)\|_{L^{2}(0,1)}+\|\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\big).$
+Taking into account (3) and (3.6), we obtain
+$\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad\leq C\exp{\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}}$
+$\qquad\quad\times\Big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\Big)$
+$\qquad\quad\times\big(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\big).$ (3.8)
+Next, since $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, we use the triangle inequality as follows:
+$\|u_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}\leq\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\|\psi_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))},$
+$\sqrt{\alpha}\|\partial_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))},$
+$\|\sqrt{q_{+,\varepsilon}}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
+Note that $\psi_{\varepsilon}(t,x)\in\operatorname{span}\{1,x\}\cdot g_{j,\varepsilon}(t)$, so
+$\|\psi_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad\leq C(T)\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big).$
+Combining this with (3) yields (3.1), which completes the proof of Proposition 3.1. ∎
+Remark 3.2.
+The last term in (3.1) corresponds to the explicit contribution of the lifting function $\psi_{\varepsilon}$; since $\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}$ is $L^{\infty}$-moderate and $g_{0,\varepsilon},g_{1,\varepsilon}$ are $H^{1}$-moderate, this contribution is moderate as well.
+Proposition 3.2.
+Let $u_{\varepsilon}$ solve the regularized non-homogeneous Dirichlet problem (3.2) with data
+$f_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1)),\quad u_{0,\varepsilon}\in H^{1}_{0}(0,1),\quad g_{0,\varepsilon},\,g_{1,\varepsilon}\in H^{1}([0,T]).$
+Assume the coefficients satisfy
+$a_{\varepsilon},\,\partial_{t}a_{\varepsilon},\,\partial_{x}a_{\varepsilon},\,b_{\varepsilon},\,\partial_{x}b_{\varepsilon}\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a_{\varepsilon}(t,x)\geq\alpha>0,$
+$q_{\varepsilon}\in L^{\infty}(0,1),\qquad q_{\varepsilon}=q_{+,\varepsilon}-q_{-,\varepsilon},\quad q_{+,\varepsilon}\geq 0,\quad q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}.$
+Then $\partial_{t}u_{\varepsilon},\,\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})\in L^{2}([0,T];L^{2}(0,1))$, and the estimate
+$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
+$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
+$\qquad\times\Big(\sqrt{\alpha}\|u_{0,\varepsilon}\|_{H^{1}_{0}(0,1)}+\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
+$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big),$ (3.9)
+where $C>0$ depends only on $T$ and the interval $(0,1)$.
+Proof.
+We decompose $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, where $\psi_{\varepsilon}$ is the boundary lifting and $w_{\varepsilon}$ satisfies homogeneous Dirichlet conditions.
+Then $w_{\varepsilon}$ solves the homogeneous problem (3.4) with source term
+$\widetilde{f}_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$
+and initial datum $u_{0,\varepsilon}-\psi_{\varepsilon}(0)\in H^{1}_{0}(0,1)$. Hence Lemma 2.4 applies and yields
+$\partial_{t}w_{\varepsilon},\ \partial_{x}(a_{\varepsilon}\partial_{x}w_{\varepsilon})\in L^{2}(0,T;L^{2}(0,1)).$
+Using the triangle inequality and the explicit form of $\psi_{\varepsilon}$, we obtain the corresponding bounds for $u_{\varepsilon}$.
+Moreover, since $\partial_{x}u_{\varepsilon}=\partial_{x}w_{\varepsilon}+\partial_{x}\psi_{\varepsilon}$ and $\partial_{x}\psi_{\varepsilon}\in L^{2}(0,T)$, we have $\partial_{x}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$.
+Finally,
+$a_{\varepsilon}\partial_{x}^{2}u_{\varepsilon}=\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})-(\partial_{x}a_{\varepsilon})\partial_{x}u_{\varepsilon},$
+and $a_{\varepsilon}\geq\alpha>0$, $\partial_{x}a_{\varepsilon}\in L^{\infty}$ imply $\partial_{x}^{2}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$. This yields (3.2). ∎
 Theorem 3.3 (Existence of very weak solutions).
-Let the coefficients $q\in\mathcal{D}^{\prime}(0,1)$, $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, and the boundary data $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$.
-Then, the initial-boundary value problem (4.2) admits a very weak solution in the sense of Definition 6.
+Let
+$a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\quad q\in\mathcal{D}^{\prime}(0,1),\quad f\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\quad u_{0}\in\mathcal{D}^{\prime}(0,1),$
+and let the boundary data satisfy $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. Assume that there exist regularisations
+$(a_{\varepsilon}),\ (b_{\varepsilon}),\ (q_{\varepsilon}),\ (f_{\varepsilon}),\ (u_{0,\varepsilon}),\ (g_{0,\varepsilon}),\ (g_{1,\varepsilon})$
+as in Definition 6, with $a_{\varepsilon}(t,x)\geq\alpha>0$ uniformly in $\varepsilon$, and such that $q_{\varepsilon}$ is log-$L^{\infty}$-moderate in the sense that
+$\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\leq C|\log\varepsilon|.$
+Then the problem (3.1) admits a very weak solution in the sense of Definition 6.
 Proof.
-Since equation (4.2) involves products of distributions, we replace it with a regularised problem using smooth approximations of all data and coefficients.
-Let $(q_{\varepsilon})$ be an $L^{\infty}$-moderate regularisation of $q$, $(a_{\varepsilon})$ a $W^{1,\infty}$-moderate regularisation of $a$, with $a_{\varepsilon}\geq\alpha>0$, $(b_{\varepsilon})$ a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation of $b$, $(u_{0,\varepsilon})$ an $H^{1}$-moderate regularisation of $u_{0}$, $(g_{0,\varepsilon}),(g_{1,\varepsilon})$ time-distributional moderate regularisations of $g_{0},g_{1}$, and $(f_{\varepsilon})$ an $L^{2}$-moderate regularisation of $f$.
-Then, for each $\varepsilon\in(0,1]$, consider the regularised problem (4.3).
-For each fixed $\varepsilon\in(0,1]$, the coefficients and data in (4.3) are smooth, and the diffusion coefficient $a_{\varepsilon}$ is uniformly elliptic.
-Hence, by classical parabolic theory and the energy estimate (see Corollary 3.2), there exists a unique weak solution
-$u_{\varepsilon}\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1)).$
-Moreover, Corollary 3.2 implies the energy estimate
-$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
-$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\qquad\times\left(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\right).$ (4.4)
-Using the moderateness assumptions for all regularised nets, the right-hand side is bounded by $C\varepsilon^{-N}$ for some $N\in\mathbb{N}$, uniformly in $\varepsilon\in(0,1]$. Thus,
-$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},$
-so $(u_{\varepsilon})$ is moderate in $H^{1}L^{2}\cap L^{2}H^{2}$, and therefore, by Definition 6, $(u_{\varepsilon})$ is a very weak solution to (4.2). ∎
+Since (3.1) involves products of distributions, we consider the regularised problem (3.2) for each $\varepsilon\in(0,1]$, where $a_{\varepsilon},b_{\varepsilon},q_{\varepsilon},f_{\varepsilon},u_{0,\varepsilon},g_{0,\varepsilon},g_{1,\varepsilon}$ are smooth regularisations satisfying the moderateness assumptions of Definition 6 and $a_{\varepsilon}\geq\alpha>0$.
+For fixed $\varepsilon$, the coefficients are bounded and $a_{\varepsilon}$ is uniformly elliptic, hence there exists a unique weak solution $u_{\varepsilon}$ to (3.2).
+Moreover, Proposition 3.2 yields
+$\partial_{t}u_{\varepsilon},\ \partial_{x}^{2}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$
+together with the estimate (3.2).
+We now show moderateness of $(u_{\varepsilon})$.
+By the assumptions on the regularisations, we have polynomial bounds for $\|u_{0,\varepsilon}\|_{H^{1}_{0}}$, $\|f_{\varepsilon}\|_{L^{2}}$, $\|a_{\varepsilon}\|_{W^{1,\infty}}$, $\|\partial_{t}a_{\varepsilon}\|_{L^{\infty}}$, $\|b_{\varepsilon}\|_{L^{\infty}}$, and log-bounds for $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}}$ and $\|q_{-,\varepsilon}\|_{L^{\infty}}$.
+Consequently,
+$\exp\!\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}}+2\|q_{-,\varepsilon}\|_{L^{\infty}}\right)\right\}\leq C\,\varepsilon^{-M}$
+for some $M\in\mathbb{N}$, and all remaining factors on the right-hand side of (3.2) are bounded by $C\varepsilon^{-M}$ as well. Hence there exist $C>0$ and $N\in\mathbb{N}$ such that
+$\|\partial_{t}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\leq C\,\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
+This is exactly the moderateness condition in Definition 6. Therefore $(u_{\varepsilon})$ is a very weak solution to (3.1). ∎
 Definition 7 (Negligibility).
-Let $(u_{\varepsilon})$, $(\tilde{u}_{\varepsilon})$ be two nets in $L^{\infty}([0,T];L^{2}(0,1))$.
-We say that the net $(u_{\varepsilon}-\tilde{u}_{\varepsilon})$ is $L^{2}$-negligible if, for every $N\in\mathbb{N}$ there exist a constant $C>0$ such that
-$\|u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)\|_{L^{2}}\leq C\varepsilon^{N},$
+Let $(u_{\varepsilon})$, $(\widetilde{u}_{\varepsilon})$ be two nets in $L^{\infty}([0,T];L^{2}(0,1))$.
+We say that the net $(u_{\varepsilon}-\widetilde{u}_{\varepsilon})$ is $L^{\infty}(L^{2})$-negligible if, for every $N\in\mathbb{N}$ there exist a constant $C>0$ such that
+$\|u_{\varepsilon}(t,\cdot)-\widetilde{u}_{\varepsilon}(t,\cdot)\|_{L^{2}}\leq C\varepsilon^{N},$
 for all $\varepsilon\in(0,1]$ uniformly in $t\in[0,T]$. The constant $C$ may depends on $N$ but not on $\varepsilon$.
@@
 and
-$\left\{\begin{array}[]{l}\partial_{t}\tilde{u}_{\varepsilon}(t,x)-\partial_{x}(\tilde{a}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon})(t,x)+\tilde{b}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon}(t,x)+\tilde{q}_{\varepsilon}(x)\tilde{u}_{\varepsilon}(t,x)\\ \hfill=\tilde{f}_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ \tilde{u}_{\varepsilon}(0,x)=\tilde{u}_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ \tilde{u}_{\varepsilon}(t,0)=\tilde{g}_{0,\varepsilon}(t),\quad\tilde{u}_{\varepsilon}(t,1)=\tilde{g}_{1,\varepsilon}(t),\quad t\in[0,T].\end{array}\right$.
+$\left\{\begin{array}[]{l}\partial_{t}\widetilde{u}_{\varepsilon}(t,x)-\partial_{x}(\widetilde{a}_{\varepsilon}(t,x)\partial_{x}\widetilde{u}_{\varepsilon})(t,x)+\widetilde{b}_{\varepsilon}(t,x)\partial_{x}\widetilde{u}_{\varepsilon}(t,x)+\widetilde{q}_{\varepsilon}(x)\widetilde{u}_{\varepsilon}(t,x)\\ \hfill=\widetilde{f}_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ \widetilde{u}_{\varepsilon}(0,x)=\widetilde{u}_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ \widetilde{u}_{\varepsilon}(t,0)=\widetilde{g}_{0,\varepsilon}(t),\quad\widetilde{u}_{\varepsilon}(t,1)=\widetilde{g}_{1,\varepsilon}(t),\quad t\in[0,T].\end{array}\right$.
 (3.11)
-Definition 8 (Uniqueness of the very weak solution).
-Let $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, $g_{0},\,g_{1}\in\mathcal{D}^{\prime}([0,T])$.
-We say that the initial/boundary problem (4.2) has a unique very weak solution, if for all $W^{1,\infty}(L^{2})$, log-$W^{1,\infty}$-moderate nets $a_{\varepsilon}$, $\tilde{a}_{\varepsilon}$, $b_{\varepsilon}$, $\tilde{b_{\varepsilon}}$, such that $(a_{\varepsilon}-\tilde{a}_{\varepsilon})$, $(b_{\varepsilon}-\tilde{b}_{\varepsilon})$ are $W^{1,\infty}$-negligible and log-$L^{\infty}(W^{1,\infty})$-negligible; $L^{\infty}$-moderate nets $q_{\varepsilon}$, $\tilde{q}_{\varepsilon}$, such that $(q_{\varepsilon}-\tilde{q}_{\varepsilon})$ are $L^{\infty}$-negligible; time-distributional moderate nets $g_{0,\varepsilon},\,g_{1,\varepsilon},\,\tilde{g}_{0,\varepsilon},\,\tilde{g}_{1,\varepsilon}$, such that $(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})$, $(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})$ are $H^{1}$-negligible; for all $L^{2}$-moderate nets $f_{\varepsilon},\,\tilde{f}_{\varepsilon}$, such that $(f_{\varepsilon}-\tilde{f}_{\varepsilon})$ are $L^{2}$-negligible; for all $H^{1}$-moderate nets $u_{0,\varepsilon},\,\tilde{u}_{0,\varepsilon}$ such that $(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})$ is $H^{1}$-negligible, we have that $u_{\varepsilon}-\tilde{u}_{\varepsilon}$ is $H^{1}(L^{2})\cap L^{2}(H^{2})$-negligible.
+Definition 8 (Uniqueness of very weak solutions).
+We say that the initial–boundary value problem (3.1) has a unique very weak solution if the following holds:
+Let $(a_{\varepsilon},b_{\varepsilon},q_{\varepsilon},u_{0,\varepsilon},g_{0,\varepsilon},g_{1,\varepsilon},f_{\varepsilon})$ and $(\widetilde{a}_{\varepsilon},\widetilde{b}_{\varepsilon},\widetilde{q}_{\varepsilon},\widetilde{u}_{0,\varepsilon},\widetilde{g}_{0,\varepsilon},\widetilde{g}_{1,\varepsilon},\widetilde{f}_{\varepsilon})$ be two families of regularisations satisfying the assumptions of Definition 6, and let $u_{\varepsilon}$, $\widetilde{u}_{\varepsilon}$ be the corresponding solutions of (3.10) and (3.11).
+Assume in addition that the differences of the regularisations are negligible in the corresponding norms, i.e.
+$a_{\varepsilon}-\widetilde{a}_{\varepsilon}\ \text{is $W^{1,\infty}$-negligible},\qquad b_{\varepsilon}-\widetilde{b}_{\varepsilon}\ \text{is log-$L^{\infty}(W^{1,\infty})$-negligible},$
+$q_{\varepsilon}-\widetilde{q}_{\varepsilon}\ \text{is $L^{\infty}$-negligible},\qquad u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}\ \text{is $H^{1}$-negligible},$
+$g_{j,\varepsilon}-\widetilde{g}_{j,\varepsilon}\ \text{is $H^{1}(0,T)$-negligible for }j=0,1,\qquad f_{\varepsilon}-\widetilde{f}_{\varepsilon}\ \text{is $L^{2}(0,T;L^{2})$-negligible}.$
+Then
+$u_{\varepsilon}-\widetilde{u}_{\varepsilon}\ \text{is $H^{1}(0,T;L^{2}(0,1))\cap L^{2}(0,T;H^{2}(0,1))$-negligible}.$
 Theorem 3.4 (Uniqueness of the very weak solution).
@@
 Proof.
-Let us denote by $u_{\varepsilon}$ and $\tilde{u}_{\varepsilon}$ the families of solutions for the initial/boundary problems (4.5) and (4.6) correspondingly.
-Let $U_{\varepsilon}$ represents the difference between these nets $U_{\varepsilon}(t,\cdot):=u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)$ then $U_{\varepsilon}$ solves the equation
-$\left\{\begin{array}[]{l}\partial_{t}U_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}U_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}U_{\varepsilon}(t,x)+q_{\varepsilon}(x)U_{\varepsilon}(t,x)\\ \hfill=F_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ U_{\varepsilon}(0,x)=(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})(x),\,\,\,x\in(0,1),\\ U_{\varepsilon}(t,0)=(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})(t),\,U_{\varepsilon}(t,1)=(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})(t),\,\,t\in[0,T],\end{array}\right$.
+Let us denote by $u_{\varepsilon}$ and $\widetilde{u}_{\varepsilon}$ the families of solutions for the initial/boundary problems (3.10) and (3.11) correspondingly.
+Let $U_{\varepsilon}$ represents the difference between these nets $U_{\varepsilon}(t,\cdot):=u_{\varepsilon}(t,\cdot)-\widetilde{u}_{\varepsilon}(t,\cdot)$ then $U_{\varepsilon}$ solves the equation
+$\left\{\begin{array}[]{l}\partial_{t}U_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}U_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}U_{\varepsilon}(t,x)+q_{\varepsilon}(x)U_{\varepsilon}(t,x)\\ \hfill=F_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ U_{\varepsilon}(0,x)=(u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon})(x),\,\,\,x\in(0,1),\\ U_{\varepsilon}(t,0)=(g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon})(t),\,U_{\varepsilon}(t,1)=(g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon})(t),\,\,t\in[0,T],\end{array}\right$.
 (3.12)
 where we set
-$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\left(\big(\tilde{a}_{\varepsilon}(t,x)-a_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}\right)(t,x)+\big(\tilde{b}_{\varepsilon}(t,x)-b_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}(t,x)$
-$+\big(\tilde{q}_{\varepsilon}(x)-q_{\varepsilon}(x)\big)\tilde{u}_{\varepsilon}(t,x)+f_{\varepsilon}(t,x)-\tilde{f}_{\varepsilon}(t,x)$
+$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\left(\big(\widetilde{a}_{\varepsilon}(t,x)-a_{\varepsilon}(t,x)\big)\partial_{x}\widetilde{u}_{\varepsilon}\right)(t,x)+\big(\widetilde{b}_{\varepsilon}(t,x)-b_{\varepsilon}(t,x)\big)\partial_{x}\widetilde{u}_{\varepsilon}(t,x)$
+$+\big(\widetilde{q}_{\varepsilon}(x)-q_{\varepsilon}(x)\big)\widetilde{u}_{\varepsilon}(t,x)+f_{\varepsilon}(t,x)-\widetilde{f}_{\varepsilon}(t,x)$
 for the source term to the initial/boundary problem (3.12).
-Taking the $H^{1}(L^{2})\cap L^{2}(H^{2})$-norm of the $U_{\varepsilon}$ by using (4), we obtain
+Taking the $H^{1}(L^{2})\cap L^{2}(H^{2})$-norm of the $U_{\varepsilon}$ by using (3.2), we obtain
 $\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
-$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
-$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\quad\times\left(\|U_{\varepsilon}(0,\cdot)\|_{L^{2}}+\|F_{\varepsilon}\|_{L^{2}}+\|U_{\varepsilon}(\cdot,0)\|_{H^{1}}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}}\right)\Bigg).$
+$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
+$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
+$\qquad\times\Big(\sqrt{\alpha}\|U_{\varepsilon}(0,\cdot)\|_{H^{1}_{0}(0,1)}+\|F_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|U_{\varepsilon}(\cdot,0)\|_{H^{1}([0,T])}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}([0,T])}\Big)$
+$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|U_{\varepsilon}(\cdot,0)\|_{H^{1}([0,T])}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}([0,T])}\big)\Big).$
 For the source term $F_{\varepsilon}$, we have
-$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\partial_{x}((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}$
-$+\|(\tilde{b}_{\varepsilon}-b_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$+\|(\tilde{q}_{\varepsilon}-q_{\varepsilon})\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$ (4.8)
+$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\partial_{x}((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}$
+$+\|(\widetilde{b}_{\varepsilon}-b_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$+\|(\widetilde{q}_{\varepsilon}-q_{\varepsilon})\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$ (3.13)
 For the first term in the right-hand side we have
-$\partial_{x}\big((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\big)=\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}+(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}^{2}\tilde{u}_{\varepsilon}.$
+$\partial_{x}\big((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}\big)=\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}+(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}^{2}\widetilde{u}_{\varepsilon}.$
 Hence
-$\|\partial_{x}\cdot((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|$ $_{L^{2}([0,T];L^{2}(0,1))}\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
-By Proposition 2.4 and Corollary 3.2, this quantity is finite and moderate. Then, we have
-$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
+$\|\partial_{x}\cdot((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon})\|$ $_{L^{2}([0,T];L^{2}(0,1))}\leq\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\quad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
+By Lemma 2.4 and Proposition 3.2, this quantity is finite and moderate. Then, we have
+$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\quad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$+\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$+\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
 Combining the initial condition of (3.12), we obtain
 $\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
-$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
-$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\qquad\Big(\|u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
-Considering the negligibility of the nets $\tilde{a}_{\varepsilon}-a_{\varepsilon}$, $\tilde{b}_{\varepsilon}-b_{\varepsilon}$, $q_{\varepsilon}-\tilde{q}_{\varepsilon}$, $u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}$, $g_{\varepsilon}-\tilde{g}_{\varepsilon}$, $f_{\varepsilon}-\tilde{f}_{\varepsilon}$, we get
-$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
-$\qquad\leq\frac{C_{0}}{\alpha}\Bigg(\varepsilon^{-N_{0}}\Big(C_{1}\varepsilon^{-N_{1}}+C_{2}\varepsilon^{-N_{2}}+\frac{1}{\sqrt{\alpha}}\left(C_{3}\varepsilon^{-N_{3}}+C_{4}\varepsilon^{-N_{4}}+C_{5}\varepsilon^{-N_{5}}\right)\Big)$
-$\qquad\times\Big(1+\frac{1}{\sqrt{\alpha}}\left(C_{6}\varepsilon^{-N_{6}}+C_{7}\varepsilon^{-N_{7}}\right)+C_{8}\varepsilon^{-N_{8}}\Big)\big(C_{9}\varepsilon^{N_{9}}+C_{10}\varepsilon^{N_{10}}\varepsilon^{-N_{11}}$
-$\qquad+C_{11}\varepsilon^{N_{12}}\varepsilon^{-N_{13}}+C_{12}\varepsilon^{N_{14}}\varepsilon^{-N_{15}}+C_{13}\varepsilon^{N_{16}}\varepsilon^{-N_{17}}+C_{14}\varepsilon^{N_{18}}+C_{15}\varepsilon^{N_{19}}+C_{16}\varepsilon^{N_{20}}\big)\Bigg)$
-for some $C_{j}>0$, j=0,,16j=0,...,16, $N_{k}\in\mathbb{N}$, $k=0,\ldots,8,\,11,\,13,\,15,\,17$, and all
-$N_{9},\,N_{10},\,N_{12},\,N_{14},\,N_{16},\,N_{18},\,N_{19},\,N_{20}\in\mathbb{N}$. Then, for some $C_{D}>0$ and all $K\in\mathbb{N}$
-$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\leq C_{D}\varepsilon^{K},$
-and this completes the proof of Theorem 4.2. ∎
+$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
+$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
+$\qquad\times\Big(\|u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
+$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big).$
+By Proposition 3.2 (applied to $\widetilde{u}_{\varepsilon}$) and the moderateness assumptions, the nets
+$\|\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})},\qquad\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})},\qquad\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})}$
+are moderate. On the other hand, by the negligibility assumptions in Definition 8, the nets
+$\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}W^{1,\infty}},\quad\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}W^{1,\infty}},\quad\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}},\quad\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}(0,T;L^{2})}$
+are negligible. Hence, each term on the right-hand side of (3) is negligible, and therefore
+$\|F_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\ \text{is negligible.}$
+Similarly, $U_{\varepsilon}(0,\cdot)=u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}$ is $H^{1}$-negligible and the boundary differences $U_{\varepsilon}(\cdot,0)$ and $U_{\varepsilon}(\cdot,1)$ are $H^{1}(0,T)$-negligible.
+Applying estimate (3.2) to (3.12) yields that
+$\|U_{\varepsilon}\|_{H^{1}(0,T;L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}(0,T;H^{2}(0,1))}$
+is negligible. This proves uniqueness of the very weak solution in the sense of Definition 8. ∎
 Theorem 3.5 (Consistency).
 Assume that $a\in W^{1,\infty}([0,T];W^{1,\infty}(0,1))$,
-$b\in L^{\infty}([0,T];W^{1,\infty}(0,1))$, $q\in L^{\infty}(0,1)$, $g_{0},\,g_{1}\in H^{1}([0,T])$, and let $a_{\varepsilon}$ be any $W^{1,\infty}$-regularisation of $a$, $b_{\varepsilon}$ be any log-$L^{\infty}(W^{1,\infty})$-regularisation of $b$, $(q_{\varepsilon})$ be any $L^{\infty}$-regularisation of $q$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$ are any time-distributional regularisations of $g_{0}$ and $g_{1}$ respectively, that is
+$b\in L^{\infty}([0,T];W^{1,\infty}(0,1))$, $q\in L^{\infty}(0,1)$, $g_{0},\,g_{1}\in H^{1}([0,T])$.
+Assume moreover that $u_{0}\in H^{1}_{0}(0,1)$.
+Let $u$ be the unique classical (weak) solution of (3.1), i.e.
+$u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1)).$
+Then for any families $a_{\varepsilon}$, $b_{\varepsilon}$, $q_{\varepsilon}$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$, $u_{0,\varepsilon}$ and $f_{\varepsilon}$ such that
 $\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
-$\|q_{\varepsilon}-q\|_{L^{\infty}(0,1)}\to 0,\quad\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0$
-as $\varepsilon\to 0$.
-Let the initial condition satisfies $u_{0}\in H^{1}(0,1)$.
-Let $u$ be a very weak solution of the initial/boundary problem (4.2).
-Then for any families $a_{\varepsilon}$, $b_{\varepsilon}$, $q_{\varepsilon}$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$, $u_{0,\varepsilon}$ and $f_{\varepsilon}$ such that
+$\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\to 0,\quad\|u_{0}-u_{0,\varepsilon}\|_{H^{1}_{0}(0,1)}\to 0,$
+$\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
+as $\varepsilon\to 0$, any representative $(u_{\varepsilon})$ of the very weak solution converges to $u$ in the sense that
+$\|u-u_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|u-u_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0,\qquad\varepsilon\to 0.$
+Proof.
+Let us denote $\widetilde{U}_{\varepsilon}$ as the difference between $u$ and $u_{\varepsilon}$, that is
+$\widetilde{U}_{\varepsilon}(t,x):=u(t,x)-u_{\varepsilon}(t,x).$
+Then the net $\widetilde{U}_{\varepsilon}$ is a solution to the initial/boundary problem
+$\left\{\begin{array}[]{l}\partial_{t}\widetilde{U}_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}\widetilde{U}_{\varepsilon})(t,x)+b_{\varepsilon}\partial_{x}\widetilde{U}_{\varepsilon}+q_{\varepsilon}(x)\widetilde{U}_{\varepsilon}(t,x)=F_{\varepsilon}(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ \widetilde{U}_{\varepsilon}(0,x)=(u_{0}-u_{0,\varepsilon})(x),\quad x\in(0,1),\\ \widetilde{U}_{\varepsilon}(t,0)=(g_{0}-g_{0,\varepsilon})(t),\quad\widetilde{U}_{\varepsilon}(t,1)=(g_{1}-g_{1,\varepsilon})(t),\quad t\in[0,T],\end{array}\right$.
+(3.14)
+where
+$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\cdot\left(\big(a_{\varepsilon}(t,x)-a(t,x)\big)\partial_{x}u\right)(t,x)+\big(b_{\varepsilon}(t,x)-b(t,x)\big)\cdot\partial_{x}u(t,x)$
+$+\big(q_{\varepsilon}(x)-q(x)\big)u(t,x)+f(t,x)-f_{\varepsilon}(t,x).$
+Since $u\in L^{2}([0,T];H^{2}(0,1))$, each term in $F_{\varepsilon}$ tends to zero in $L^{2}([0,T];L^{2}(0,1))$, hence $\|F_{\varepsilon}\|_{L^{2}}\to 0$.
+Analogously to Theorem 3.4 we have that
+$\|\widetilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\widetilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
+$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
+$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
+$\qquad\times\Big(\|u_{0}-{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|a-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|\partial_{x}(a-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|b-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|{u}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f-{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
+$\qquad+\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
+$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big).$
+Since
 $\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
@@
 $\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
-as $\varepsilon\to 0$, any representative $(u_{\varepsilon})$ of the very weak solution converges as
-$\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{H^{1}([0,T];L^{2}(0,1))}+\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
-for $\varepsilon\to 0$ to the unique classical solution $u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1))$ of the initial/boundary problem (4.2) given by Theorem 3.1.
-Proof.
-Let us denote $\tilde{U}_{\varepsilon}$ as the difference between $u$ and $u_{\varepsilon}$, that is
-$\tilde{U}_{\varepsilon}(t,x):=u(t,x)-u_{\varepsilon}(t,x).$
-Then the net $\tilde{U}_{\varepsilon}$ is a solution to the initial/boundary problem
-$\left\{\begin{array}[]{l}\partial_{t}\tilde{U}_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon})(t,x)+b_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon}+q_{\varepsilon}(x)\tilde{U}_{\varepsilon}(t,x)=F_{\varepsilon}(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ \tilde{U}_{\varepsilon}(0,x)=(u_{0}-u_{0,\varepsilon})(x),\quad x\in(0,1),\\ \tilde{U}_{\varepsilon}(t,0)=(g_{0}-g_{0,\varepsilon})(t),\quad\tilde{U}_{\varepsilon}(t,1)=(g_{1}-g_{1,\varepsilon})(t),\quad t\in[0,T],\end{array}\right$.
-(4.9)
-where
-$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\cdot\left(\big(a_{\varepsilon}(t,x)-a(t,x)\big)\partial_{x}u\right)(t,x)+\big(b_{\varepsilon}(t,x)-b(t,x)\big)\cdot\partial_{x}u(t,x)$
-$+\big(q_{\varepsilon}(x)-q(x)\big)u(t,x)+f(t,x)-f_{\varepsilon}(t,x)$
-Analogously to Theorem 4.2 we have that
-$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
-$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
-$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
-[…]
-$\qquad\times\Big(\|u_{0}-{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|{a}_{\varepsilon}-a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|\partial_{x}({a}_{\varepsilon}-a)\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|{b}_{\varepsilon}-b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|{q}_{\varepsilon}-q\|_{L^{\infty}(0,1)}\|{u}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f-{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
-$\qquad+\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
-Since
-$\|a_{\varepsilon}-a\|_{W^{1,\infty}([0,T];W^{1,\infty}(0,1))}\to 0,\quad\|b_{\varepsilon}-b\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\to 0,$
-$\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\to 0,\quad\|u_{0}-u_{0,\varepsilon}\|_{H^{1}(0,1)}\to 0,$
-$\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])},\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])},\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
-for $\varepsilon\to 0$ and $u$ is a very weak solution of the initial/boundary problem (4.2), we get
-$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
-for $\varepsilon\to 0$. This proves Theorem 4.3.
+for $\varepsilon\to 0$ and $u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1))$, we obtain
+$\|\widetilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\widetilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0,$
+which proves Theorem 3.5. ∎
 4. Numerical Experiments